### v2

# Gauge fields in the presence of the electroweak bubble wall

Abstract
The gauge field theory of the standard electroweak model in the presence of the electroweak bubble wall is investigated with an eye toward its applications to microscopic phenomena, which are supposed to have occurred during the phase transition in the early universe. The asymptotic fields are defined anew so that the effects of the position-dependent Higgs condensate are taken into account through the position-dependent $W$ and $Z$ boson masses. A novel method of massive gauge field quantization in the $R_{\xi}$-gauge with $\xi=1$ is proposed for the case of the position-dependent masses. Our procedure is based on the eigenfunction expansion method associated with second-order differential operators, i.e., a sort of generalized Fourier expansion.The commutation relations of creation and annihilation operators of various wave propagation modes are given in terms of the so-called spectral function. The decoupling of unphysical states from the physical S-matrix is also investigated along the line of Kugo-Ojima’s quartet mechanism on the basis of the BRST symmetry. It is pointed out that one of the quartet fields is not merely the unphysical scalar field but should be a linear combination of the unphysical scalar and the gauge fields. The physical and unphysical polarizations of the gauge field waves are unambiguously distinguished and this will help us evaluate the friction caused by the physical polarization states of $W$ and $Z$ boson waves on the bubble wall during the phase transition in the early universe.
1 Introduction
The present paper is a sequel to the author’s pervious one kubota1 , in which two-point Green’s functions of scalar, spinor and vector fields in the presence of the electroweak bubble wall were constructed. The purpose of Ref.kubota1 and of the present paper is to prepare field theoretical tools for analyzing phenomena that must have occurred near the electroweak bubble wall during the expansion of the Higgs condensate in the early universe. Detailed analyses of those phenomena are indispensable for precise microscopic understanding of the magnitude of the friction caused on the bubble wall by the plasma consisting of standard model particles. The eventual goal of these considerations is to evaluate hopefully the bubble wall velocity as accurately as possible and to get information on the gravitational wave emitted during the bubble expansion.
The standard model phase transition is smooth cross-over, but could be of first-order if the electroweak model is extended somehow in the future kajantie - grojean . The bubble wall velocity during the phase transition in the early universe is one of the key parameters to determine the strength and spectral shape of the gravitational wave that was presumably emitted during the first-order phase transition witten - kosowsky2 . The gravitational wave projects planned in the near future such as LISA caprini1 - auclair , DECIGO kawamura0 - kawamura , Taiji ruan and TianQin luo - luo2 are expected to have suitable sensitivity to probe the electroweak phase transition and therefore detailed analysis on the bubble wall velocity is pressingly important. It is also of great significance for the electroweak baryogenesis kuzmin - garbrecht .
The Higgs vacuum expectation value varies through the bubble wall interface and therefore the standard model particles have position dependent masses. The Green’s functions in such a case are expressed in kubota1 in the form of integral representations, where the so-called eigenfunction expansion method initiated by Weyl weyl , developed further by Stone stone and completed in its final form by Titchmarsh titchmarsh and Kodaira kodaira1 - kodaira2 was shown to be an indispensable and at the same time the most suitable technique. (See also yoshida - yosida2 for a pedagogical exposition.) It has been shown in kubota1 that the spectral function of the eigenvalue problem of a self-adjoint differential operator plays a crucial role in constructing the Green’s functions.
The Green’s functions can be obtained simply by solving partial differential equations under given boundary conditions without touching upon the procedure of field quantization or the details of the Hilbert space. From the standpoint of quantum field theory, however, the method of field quantization must be reexamined before applying the Green’s functions to quantum phenomena. Since masses are position dependent, the notions of asymptotic fields and their wave equations have to be reconsidered. By studying the various solutions of the wave equations, we are able to define a set of operators which are analogous to creation and annihilation operators, but are no more associated with conventional particle pictures. They can create or annihilate only particular wave propagation modes. The electroweak bubble wall breaks both translational and Lorentz invariance, and therefore we are not able to use the conventional approach of classifying particles according to their spin and mass. Since the particle picture is veiled, the Hilbert space constructed by applying creation operators on the vacuum will become necessarily different from the usual Fock space
In the electroweak gauge theory, the Hilbert space contains several kinds of unphysical negative norm states as well as physical but zero-norm states. In the early days of non-abelian gauge theories, Kugo and Ojima kugoojima1 - kugoojima5 developed a powerful method of classifying physical and unphysical states by employing the BRST (Becchi, Rouet, Stora brst3 -brst1 and Tyutin tyutin ) invariance. They advocated the so-called “BRST quartet mechanism", by which unwanted states such as ghost, anti-ghost, scalar polarization and unphysical scalar field states conspire together to decouple from the physical S-matrix. In the presence of the electroweak bubble wall, although we have still the BRST invariance of the Lagrangian, the Lorentz invariance is no more respected and it is far from obvious how the BRST quartet mechanism works.
The purpose of the present paper is to scrutinize how and to what extent the presence of the electroweak bubble modifies the framework and notions in gauge theories, such as asymptotic fields, field quantization method,  [^1] See Refs. boedeker2 - ai3 for other approaches to field quantization under bubble wall backgrounds. creation and annihilation operators, classification of vector field polarizations, the decoupling of unphysical states from the physical S-matrix, and so forth. It should be emphasized that these are not totally academic problems but are strongly motivated by the recent intense studies on the bubble wall expansion dynamics in the early universe.
The electroweak bubble wall feels differential vacuum energy across the bubble interface and is driven forward so that the bubble expands. In addition, as argued in boedeker1 , there occur friction effects caused by plasma particles passing through the bubble wall. Particles get massive when they pass through the wall from symmetry-restored to symmetry-broken regions. The energy and the transverse momentum of the particles are conserved and some amount of momentum in the direction perpendicular to the wall necessarily decreases in correspondence with the increase of the mass. The momentum lost by plasma particles is transferred to the wall, as a result of which the wall feels friction impeding the bubble expansion.
In the meanwhile the effects due to the transition radiation ginzburgfrank - jackson have also been studied in boedeker2 - GouttennoireJinnoSala . Namely, the process of a particle impinging on the wall and emitting another particle (or several particles) was investigated. It was argued that soft $W$- and $Z$- vector boson emission dominates the friction on the wall. The pressure exerted on the wall is enhanced and is featured by logarithmic terms that come from the soft region in the phase space. The fixed order calculation of the thermal pressure breaks down quite possibly and the method of summation of multiple soft gauge boson emission effects are proposed. According to the common lore of infrared singularities in field theories, a great care must be exerted to handle the masses, in particular, of gauge bosons in infrared problems. The present paper is intended to offer theoretical bases for such subtleties connected with the gauge fields whose masses are position-dependent.
The structure of the present article is as follows. In Section 2, the electroweak bubble wall is prescribed by a differential equation ensuring the absence of tadpoles. In Section 3, the asymptotic fields for the case of position-dependent mass are defined and are expanded in terms of mode functions by taking the scalar field case as an example. The procedure of the scalar field quantization is exemplified by using the spectral function. The gauge fixing procedure, the ghost and anti-ghost fields and the BRST symmetry are summarized in Section 4 for the sake of preparation to quantize the $Z$-boson field in Section 5. The properties of auxiliary fields that are introduced through the gauge fixing procedure are given in Section 6 and it is argued that the Kugo-Ojima’s quartet mechanism works well by modifying one of the quartet fields appropriately. The physical polarization states are explicitly given in Section 7. Section 8 is devoted to the summary of the present paper.
2 The electroweak bubble wall
First of all we would like to specify the electroweak bubble wall to be treated in the present paper, lest our scope should become too much stretched. Let us begin with the scalar part of the Lagrangian in the standard electroweak theory
${\cal L}_{\rm scalar}=(\nabla^{\mu}\Phi)^{{\dagger}}\nabla_{\mu}\Phi-V(\Phi^{{\dagger}}\Phi)\>.$ (1)
We parameterize the standard model Higgs doublet field $\Phi$ in the following way,
[…]
where $H$ and $\chi^{a}$ $(a=1,2,3)$ are the Higgs boson and unphysical scalar fields, respectively. The Pauli matrices are denoted by $\tau^{a}$ $(a=1,2,3)$. The vacuum expectation value $v$ is not necessarily a constant in the present context, but may depend on the space-time coordinates. We will go as far as possible without specifying the explicit form of the Higgs potential $V(\Phi^{{\dagger}}\Phi)$ in (1). This is because the standard electroweak theory will be possibly extended in the future so that the phase transition is of the first order and therefore the Higgs potential should be kept general enough in the form amenable to future modification.
Putting (8) into (1), we find a so-called tadpole term which is linear in the Higgs field $H$, namely,
${\cal L}_{\rm tadpole}=-H\left\{\square\>v+vV^{\>\prime}\left(\frac{v^{2}}{2}\right)\right\}\>,$ (9)
where the prime $(^{\prime})$ in $V^{\>\prime}(\Phi^{{\dagger}}\Phi)$ means the derivative of the potential with respect to $\Phi^{{\dagger}}\Phi$. The d’Alembertian operator in (9) is given as usual by
$\square=g^{\mu\nu}\frac{\partial}{\partial x^{\mu}}\frac{\partial}{\partial x^{\nu}}=\frac{\partial^{2}}{\partial t^{2}}-\frac{\partial^{2}}{\partial x^{2}}-\frac{\partial^{2}}{\partial y^{2}}-\frac{\partial^{2}}{\partial z^{2}}\>,$ (10)
where our convention of the metric is $g^{\mu\nu}={\rm diag}(1,-1,-1,-1)$. Here and hereafter we use the hybrid notations for the space-time coordinates, namely, $x^{\mu}=(t,x,y,z)$ together with $x^{\mu}=(x^{0},\>\cdots,\>x^{3})$. To ensure the stability, this tadpole term should vanish and we impose a condition that must be satisfied by $v$,
$\square\>v+vV^{\>\prime}\left(\frac{v^{2}}{2}\right)=0\>.$ (11)
The trivial constant solutions to (11) are $v=0$ and another one $v=v_{0}\neq 0$ satisfying $V^{\prime}(v_{0}^{\>2}/2)=0$, which correspond to symmetry-restored and symmetry-broken vacuum solutions, respectively. If, however, $v$ is not a constant but depends on one of the three space coordinates, say $z$, then the equation (11) for $v=v(z)$ becomes a non-trivial ordinary differential equation,
$-\frac{d^{2}v(z)}{dz^{2}}+v(z)V^{\>\prime}\left(\frac{v(z)^{2}}{2}\right)=0\>.$ (12)
The $z\to\pm\infty$ behavior of the solution to (12) should be a constant that is either $v=0$ or $v=v_{0}$ with $V^{\>\prime}(v_{0}^{\>2}/2)=0$. An illustrative shape of such a solution is given in Figure 1 as an example, which connects the symmetry-restored region ($z\to$-$\infty$) with the symmetry-broken region ($z\to+\infty$). The electroweak bubble wall is the region in-between around $z\sim 0$ where $v(z)$ is varying. See Ref.polyakov1 for an early attempt of making use of the solutions to (12) and also see Refs. ayala and farrar for wave propagation analyses through domain walls described by eq. (12).
[…]
Figure 1: A typical profile of the Higgs condensate $v(z)$, that connects the symmetry-restored region ($z\to$-$\infty$) and the symmetry-broken region ($z\to+\infty$).
In the present work, we simply assume the existence of such a bubble wall solution $v(z)$ as illustrated in Figure 1 without going into the detailed form of $V(\Phi^{{\dagger}}\Phi)$ or $v(z)$, either. All we require is the asymptotic behavior, $v(-\infty)=0$ and $v(+\infty)=v_{0}$  . We would like to focus our attention to scrutinize whether the conventional field theory technique works just as well in the presence of the $z$-dependent Higgs condensate $v(z)$. The procedure of field quantization itself becomes non-trivial, because all the masses of the standard model particles are now $z$-dependent. The $W$- and $Z$-boson masses for example are given, respectively, by
$M_{W}^{2}=M_{W}(z)^{2}=\frac{1}{4}g^{2}v(z)^{2},\hskip 28.45274ptM_{Z}^{2}=M_{Z}(z)^{2}=\frac{1}{4}\left(g^{2}+g^{\>\prime\>2}\right)v(z)^{2},$ (13)
where $g$ and $g^{\>\prime}$ are the $SU(2)_{L}$ and $U(1)_{Y}$ gauge couplings. We have to go into the details of the wave propagation modes along the $z$-direction, which is no more expressed as plane waves.
In the electroweak phase transition in the early universe, the bubble wall must have expanded very rapidly in the rest frame of electroweak plasma particles. The solution $v=v(z)$ to (12) on the other hand is time-independent, and the use of this solution indicates that we are not working in such a Lorentz frame but in the bubble wall rest frame. By setting $v=v(z)$, the bubble wall is supposed to be planar and this can be justified if we note that the length scale of the microscopic phenomena which we would like to analyze should be much smaller than the size of the Higgs condensate.
3 Wave propagation modes
3.1 Asymptotic fields
In the presence of the electroweak bubble wall, a generic class of standard model Heisenberg fields $\varphi(t,\vec{x}_{\perp},z)$ in the $R_{\>\xi}$ gauge with $\xi=1$ satisfies the following type of equations of motion (except for the electromagnetic and matter sector fields)
$\left\{\square+M_{W}(z)^{2}\right\}\varphi(t,\vec{x}_{\perp},z)$ $=$ $j_{\small W}(t,\vec{x}_{\perp},z)\>,$ (14)
$\left\{\square+M_{Z}(z)^{2}\right\}\varphi(t,\vec{x}_{\perp},z)$ $=$ $j_{\small Z}(t,\vec{x}_{\perp},z)\;,$ (15)
where $j_{\small W}(t,\vec{x}_{\perp},z)$ and $j_{\small Z}(t,\vec{x}_{\perp},z)$ represent interactions of $\varphi(t,\vec{x}_{\perp},z)$. These equations differ from the usual one in that the gauge boson masses, $M_{W}(z)$ and $M_{Z}(z)$, are position $z$-dependent. Due to these masses, there is no translational invariance in the $z$-direction, and we have to handle the $z$-coordinate in a different way from the transverse coordinates, $\vec{x}_{\perp}=(x,y,0)$. The analyses to be given below on (14) and those on (15) go in perfect parallel and we take only the latter equation (15) hereafter.
The formal solution to (15) can be expressed as yangfeldman - lsz2
$\varphi(t,\vec{x}_{\perp},z)$ $=$ $\varphi^{\rm in}(t,\vec{x}_{\perp},z)$ (16)
$\thinspace+\int dt^{\>\prime}\>d^{2}\vec{x}_{\perp}^{\>\prime}\>dz^{\>\prime}\>\Delta_{R}(t-t^{\>\prime},\vec{x}_{\perp}-\vec{x}_{\perp}^{\>\prime},z,z^{\>\prime};M_{Z}(*))j_{\small Z}(t^{\>\prime},\vec{x}_{\perp}^{\>\prime},z^{\>\prime})\>,$
$\varphi(t,\vec{x}_{\perp},z)$ $=$ $\varphi^{\rm out}(t,\vec{x}_{\perp},z)$ (17)
$\thinspace+\int dt^{\>\prime}\>d^{2}\vec{x}_{\perp}^{\>\prime}\>dz^{\>\prime}\>\Delta_{A}(t-t^{\>\prime},\vec{x}_{\perp}-\vec{x}_{\perp}^{\>\prime},z,z^{\>\prime};M_{Z}(*))j_{\small Z}(t^{\>\prime},\vec{x}_{\perp}^{\>\prime},z^{\>\prime})\>.$
We introduced retarded ($\Delta_{R}$) and advanced ($\Delta_{A}$) Green’s functions that satisfy, respectively, the following differential equations,
$\left\{\square+M_{Z}(z)^{2}\right\}\Delta_{R}(t-t^{\>\prime},\vec{x}_{\perp}-\vec{x}_{\perp}^{\>\prime},z,z^{\>\prime};M_{Z}(*))$ $=$ $\delta(t-t^{\>\prime})\delta^{2}(\vec{x}_{\perp}-\vec{x}_{\perp}^{\;\prime})\delta(z-z^{\>\prime})\>,$
$\left\{\square^{\;\prime}+M_{Z}(z^{\>\prime})^{2}\right\}\Delta_{R}(t-t^{\>\prime},\vec{x}_{\perp}-\vec{x}_{\perp}^{\>\prime},z,z^{\>\prime};M_{Z}(*))$ $=$ $\delta(t-t^{\>\prime})\delta^{2}(\vec{x}_{\perp}-\vec{x}_{\perp}^{\;\prime})\delta(z-z^{\>\prime})\>,$
and
$\left\{\square+M_{Z}(z)^{2}\right\}\Delta_{A}(t-t^{\>\prime},\vec{x}_{\perp}-\vec{x}_{\perp}^{\>\prime},z,z^{\>\prime};M_{Z}(*))$ $=$ $\delta(t-t^{\>\prime})\delta^{2}(\vec{x}_{\perp}-\vec{x}_{\perp}^{\;\prime})\delta(z-z^{\>\prime})\>,$
$\left\{\square^{\>\prime}+M_{Z}(z^{\>\prime})^{2}\right\}\Delta_{A}(t-t^{\>\prime},\vec{x}_{\perp}-\vec{x}_{\perp}^{\>\prime},z,z^{\>\prime};M_{Z}(*))$ $=$ $\delta(t-t^{\>\prime})\delta^{2}(\vec{x}_{\perp}-\vec{x}_{\perp}^{\;\prime})\delta(z-z^{\>\prime})\>.$
We used the notation of the d’Alembertian $\square^{\>\prime}$ defined by
[…]
We also made use of the notation $M_{Z}(*)$ as opposed to $M_{Z}(z)$ or $M_{Z}(z^{\>\prime})$, so that we keep in our mind that the dependence of the Green’s functions on $M_{Z}(z)$ or $M_{Z}(z^{\>\prime})$ is only indirect as we will confirm later. Since we do not discuss renormalization procedures in the present paper, we omit the renormalization factor in front of the incoming $\varphi^{\rm in}(t,\vec{x}_{\perp},z)$ and outgoing $\varphi^{\rm out}(t,\vec{x}_{\perp},z)$ fields. The retarded and advanced Green’s functions are non-vanishing for $t>t^{\>\prime}$ and $t^{\>\prime}>t$, respectively, and therefore $\varphi^{\rm in}(t,\vec{x}_{\perp},z)$ and $\varphi^{\rm out}(t,\vec{x}_{\perp},z)$ describe the asymptotic behavior of the field in the infinite past and infinite future, respectively (in the sense of so-called weak asymptotic condition), i.e.,
$\varphi(t,\vec{x}_{\perp},z)$ $\longrightarrow$ $\varphi^{\rm in}(t,\vec{x}_{\perp},z)\>,\hskip 34.14322pt(t\longrightarrow-\infty)\>,$ (21)
$\varphi(t,\vec{x}_{\perp},z)$ $\longrightarrow$ $\varphi^{\rm out}(t,\vec{x}_{\perp},z)\>,\hskip 28.45274pt(t\longrightarrow+\infty)\>.$ (22)
These are what we call asymptotic fields which satisfy the “free" field equations of motion,
$\left\{\square+M_{Z}(z)^{2}\right\}\varphi^{\rm in}(t,\vec{x}_{\perp},z)=0,\hskip 28.45274pt\left\{\square+M_{Z}(z)^{2}\right\}\varphi^{\rm out}(t,\vec{x}_{\perp},z)=0\>.$ (23)
The asymptotic field has been one of the basic concepts in the formulation of field theories à la Lehmann, Symanzik and Zimmermann (LSZ) lsz - lsz2 , but note that, in our case, the mass term in (23) is not a constant but is $z$-dependent. Very roughly speaking, the concept of asymptotic fields is considered in the situation in which interaction regions are restricted in space-time and forces between particles are switched off in the remote past and remote future. Nevertheless in our case, interactions with the background of Higgs condensate are still taken into account. Such a modification of the notion of asymptotic fields must be reexamined in detail carefully to see whether the LSZ techniques remain intact. In particular we should be concerned with the expansion formulas of interacting Heisenberg field operators with respect to our asymptotic fields haag1 - bogoliubov .
For now, we have to put our definition of the asymptotic fields on a sound basis by specifying, first of all, a mathematical method of constructing the retarded and advanced Green’s functions. Technical tools for elaborating the Green’s functions are prepared in the course of solving (23) and of getting all the wave propagation modes for given mass function $M_{Z}(z)$. Therefore we investigate the solutions to (23) first in Section 3.2 and then will come back later in Section 3.4 to the retarded and advanced Green’s functions.
3.2 The Klein-Gordon type real scalar field
Let us investigate wave-propagation modes of the real scalar field satisfying (23). Here and hereafter we drop the superscript “in" and “out" for simplicity, hoping that there would not arise any confusion, and we rewrite (23) again
$\left\{\square+M_{Z}(z)^{2}\right\}\varphi(t,\vec{x}_{\perp},z)=0\>.$ (24)
[…]
$e^{-iEt}\>e^{i\vec{p}_{\perp}\cdot\vec{x}_{\perp}}\>\phi(z)\>,$ (25)
where $\vec{p}_{\perp}=(p_{x},p_{y},0)$ is the wave vector in the transverse direction. By putting (25) into (24), we get an equation for $\phi(z)$, which turns out to be of the Schrödinger type
$\left\{-\frac{d^{2}}{dz^{2}}+M_{Z}(z)^{2}\right\}\phi(z;\lambda)=\lambda\phi(z;\lambda),\quad\lambda=E^{2}-|\vec{p}_{\perp}|^{2}\>,\quad(-\infty<z<+\infty)\>.$ (26)
Since this is a second order differential equation, there are two independent solutions which are denoted by $\phi_{i}(z;\lambda)\>\>(i=1,2)$ and are characterized by the initial conditions at $z=0$,
$\phi_{1}(0;\lambda)$ $=$ $1,\hskip 28.45274pt\phi_{1}^{\>\prime}(0;\lambda)=0\>,$ (27)
$\phi_{2}(0;\lambda)$ $=$ $0,\hskip 28.45274pt\phi_{2}^{\>\prime}(0;\lambda)=1\>.$ (28)
Here the prime () means the derivative with respect to $z$. Since there appear only real numbers on the right hand side of (27) and (28), the solutions $\phi_{i}(z;\lambda),\>(i=1,2)$ are both real functions.
According to Refs. weyl and stone , there exists a $2\times 2$ matrix function $\rho_{ij}(\lambda)\>\>(i,j=1,2)$ that describes the spectra of the self-adjoint differential operator on the left hand side of (26) and satisfies the completeness relation
$\int\sum_{i,j=1,2}\phi_{i}(z;\lambda)d\rho_{ij}(\lambda)\phi_{j}(z^{\prime};\lambda)=\delta(z-z^{\prime})\>.$ (29)
Here the integration $d\rho_{ij}(\lambda)$ in (29) is to be understood as the Stieltjes integral, but it turns out to be the ordinary Riemann integral if the $\lambda$-spectra in (26) do not allow discrete ones. Hereafter, for the sake of simplicity, we assume that the $\lambda$-spectra in (26) are continuous. It should be kept in our mind that the algorithm of deriving the spectral function $\rho_{ij}(\lambda)$ has been well established. See Refs. titchmarsh - yosida2 , kubota1 for more details.
3.3 Scalar field quantization in the presence of the electroweak bubble wall
Now the solution to the wave equation (24) may be expanded in terms of the complete set of the solutions as
$\varphi(t,\vec{x}_{\perp},z)$ $=$ $\int d\lambda\sum_{i=1,2}\int\frac{d^{2}\vec{p}_{\perp}}{\sqrt{(2\pi)^{2}\>2E}}$ (30)
$\times\bigg{\{}\alpha_{i}(\vec{p}_{\perp},\lambda)e^{i\vec{p}_{\perp}\cdot\vec{x}_{\perp}-iEt}+\alpha_{i}^{{\dagger}}(\vec{p}_{\perp},\lambda)e^{-i\vec{p}_{\perp}\cdot\vec{x}_{\perp}+iEt}\bigg{\}}\phi_{i}(z;\lambda)\>,$
where $E$ is subject to the “on-shell" condition and is positive
$E=\sqrt{\vec{p}^{\>2}_{\perp}+\lambda}\>>0\>.$ (31)
We should attach a superscript “in" or “out" to the coefficients $\alpha_{i}$ and $\alpha_{i}^{{\dagger}}$, but we omit it here and hereafter only for simplicity. In quantum theories the coefficients $\alpha_{i}$ and $\alpha_{i}^{{\dagger}}$ are regraded as operators satisfying certain commutation relations in such a way that the equal-time canonical commutation relations
$\left[\dot{\varphi}(t,\vec{x}_{\perp},z),\varphi(t,\vec{x}_{\perp}^{\>\prime},z^{\prime})\right]=-i\>\delta^{2}(\vec{x}_{\perp}-\vec{x}_{\perp}^{\>\prime})\delta(z-z^{\prime})\>,$ (32)
$\left[{\varphi}(t,\vec{x}_{\perp},z),\varphi(t,\vec{x}_{\perp}^{\>\prime},z^{\prime})\right]=0\>,$ (33)
$\left[\dot{\varphi}(t,\vec{x}_{\perp},z),\dot{\varphi}(t,\vec{x}_{\perp}^{\>\prime},z^{\prime})\right]=0\>,$ (34)
are realized. Here $\dot{\varphi}(t,\vec{x}_{\perp},z)$ is supposed to be canonically conjugate to $\varphi(t,\vec{x}_{\perp},z)$ and the dot ($\cdot$) means the derivative with respect to $t$. By looking at the completeness relation (29), we immediately notice that the relations (32), (33) and (34) can be derived just by postulating the following commutation relations
$\left[\alpha_{i}(\vec{p}_{\perp},\lambda),\>\alpha_{j}^{{\dagger}}(\vec{p}_{\perp}^{\>\prime},\lambda^{\prime})\right]=\delta^{2}(\vec{p}_{\perp}-\vec{p}_{\perp}^{\>\prime})\delta(\lambda-\lambda^{\prime})\frac{d\rho_{ij}(\lambda)}{d\lambda}\>,$ (35)
$\left[\alpha_{i}(\vec{p}_{\perp},\lambda),\>\alpha_{j}(\vec{p}_{\perp}^{\>\prime},\lambda^{\prime})\right]=\left[\alpha_{i}^{{\dagger}}(\vec{p}_{\perp},\lambda),\>\alpha_{j}^{{\dagger}}(\vec{p}_{\perp}^{\>\prime},\lambda^{\prime})\right]=0\>.\hskip 28.45274pt(i,j=1,2)$ (36)
These relations differ from those of creation and/or annihilation of particles by the presence of $d\rho_{ij}(\lambda)/d\lambda$ on the right hand side of (35). Note, however, that $d\rho_{ij}(\lambda)/d\lambda$ is a symmetric positive definite matrix and therefore we are able to interpret $\alpha_{i}^{{\dagger}}(\vec{p}_{\perp},\lambda)$ and $\alpha_{i}(\vec{p}_{\perp},\lambda)$ as the creation and annihilation operators, respectively, of the wave represented by (25) and labelled by $\vec{p}_{\perp}$, $\lambda=E^{2}-|\vec{p}_{\perp}|^{2}$ and i(=1i\>(=1 or $2)$. More specifically speaking, the states $\alpha_{i}^{{\dagger}}(\vec{p}_{\perp},\lambda)|0\rangle$ are all positive norm states and there is no fear of introducing negative norm states via (35) into our framework.
By employing (35) and (36), we are able to derive the four-dimensional commutation relations in the same form as in the ordinary constant mass case, i.e.,
$\left[\varphi(t,\vec{x}_{\perp},z),\>\varphi(t^{\>\prime},\vec{x}_{\perp}^{\>\prime},z^{\>\prime})\right]=i\Delta\left(t-t^{\>\prime},\vec{x}_{\perp}-\vec{x}_{\perp}^{\>\prime},z,z^{\>\prime};M_{Z}(*)\right)\>.$ (37)
Here we introduced a novel function
$\thinspace\Delta(t-t^{\>\prime},\vec{x}_{\perp}-\vec{x}_{\perp}^{\>\prime},z,z^{\>\prime};M_{Z}(*))$
$=$ $-i\int\frac{d^{2}\vec{p}_{\perp}}{(2\pi)^{2}}\>e^{i\vec{p}_{\perp}\cdot(\vec{x}_{\perp}-\vec{x}^{\;\prime}_{\perp})}\sum_{i,j=1,2}\int\phi_{i}(z;\lambda)\frac{d\rho_{ij}(\lambda)}{2E}\phi_{j}(z^{\>\prime};\lambda)\left\{e^{-iE(t-t^{\>\prime})}-e^{iE(t-t^{\>\prime})}\right\}\>,$
which is analogous to the familiar invariant $\Delta$-function. (See bjorkendrell or itzyksonzuber for example.) As will be confirmed later, when the mass-function $M_{Z}(z)$ is independent of $z$, then (LABEL:eq:deltafunction) reduces to the conventional invariant $\Delta$-function in the form of Fourier transformation. It should also be noted that the function (LABEL:eq:deltafunction) satisfies
$\left\{\square+M_{Z}(z)^{2}\right\}\Delta(t-t^{\>\prime},\vec{x}_{\perp}-\vec{x}_{\perp}^{\>\prime},z,z^{\>\prime};M_{Z}(*))=0\>,$ (39)
$\left\{\square^{\>\prime}+M_{Z}(z^{\>\prime})^{2}\right\}\Delta(t-t^{\>\prime},\vec{x}_{\perp}-{\vec{x}}_{\perp}^{\>\prime},z,z^{\>\prime};M_{Z}(*))=0\>.$ (40)
The properties at $t=t^{\>\prime}$
$\Delta(0,\vec{x}_{\perp}-{\vec{x}}_{\perp}^{\>\prime},z,z^{\>\prime};M_{Z}(*))=0\>,$ (41)
$\frac{\partial}{\partial t}\Delta(t-t^{\>\prime},\vec{x}_{\perp}-{\vec{x}}_{\perp}^{\>\prime},z,z^{\>\prime};M_{Z}(*))\bigg{|}_{t=t^{\>\prime}}=-\delta^{2}(\vec{x}_{\perp}-\vec{x}^{\>\prime}_{\perp})\delta(z-z^{\>\prime})\>,$ (42)
$\frac{\partial}{\partial t^{\;\prime}}\frac{\partial}{\partial t}\Delta(t-t^{\>\prime},\vec{x}_{\perp}-{\vec{x}}_{\perp}^{\>\prime},z,z^{\>\prime};M_{Z}(*))\bigg{|}_{t=t^{\>\prime}}=0\>.$ (43)
should also be mentioned, and these properties enable us to reach (32), (33) and (34) consistently from (37) .
3.4 The advanced and retarded Green’s functions
We are now well equipped to present the retarded and advanced Green’s functions, which are given by
$\thinspace\Delta_{R}(t-t^{\>\prime},\vec{x}_{\perp}-\vec{x}_{\perp}^{\>\prime},z,z^{\>\prime};M_{Z}(*))$
$=-\int\frac{dE\>d^{2}\vec{p}_{\perp}}{(2\pi)^{3}}e^{-iE(t-t^{\>\prime})}e^{i\vec{p}_{\perp}\cdot(\vec{x}_{\perp}-\vec{x}_{\perp}^{\>\prime})}\sum_{i,j=1,2}\int\frac{\phi_{i}(z;\lambda)d\rho_{ij}(\lambda)\phi_{j}(z^{\>\prime};\lambda)}{(E+i\varepsilon)^{2}-\vec{p}_{\perp}^{\>2}-\lambda}\>,$ (44)
$\thinspace\Delta_{A}(t-t^{\>\prime},\vec{x}_{\perp}-\vec{x}_{\perp}^{\>\prime},z,z^{\>\prime};M_{Z}(*))$
$=-\int\frac{dE\>d^{2}\vec{p}_{\perp}}{(2\pi)^{3}}e^{-iE(t-t^{\>\prime})}e^{i\vec{p}_{\perp}\cdot(\vec{x}_{\perp}-\vec{x}_{\perp}^{\>\prime})}\sum_{i,j=1,2}\int\frac{\phi_{i}(z;\lambda)d\rho_{ij}(\lambda)\phi_{j}(z^{\>\prime};\lambda)}{(E-i\varepsilon)^{2}-\vec{p}_{\perp}^{\>2}-\lambda}\>.$ (45)
It is almost obvious that (44) and (45) satisfy (LABEL:eq:retardedgreensfunction) and (LABEL:eq:advancedgreensfunction) respectively thanks to the formula (29). Using the Cauchy theorem on the complex $E$-plane by adding the integration path of an infinite semi-circle, we can confirm that $\Delta_{R}$ ($\Delta_{A}$) is non-vanishing only for $t>t^{\>\prime}$ $(t<t^{\>\prime})$. With these explicit formulas of the Green’s functions, the definition of the asymptotic fields given in Section 3.1 is well established.
3.5 The special case of $z$-independent mass (no bubble wall)
For illustration, let us discuss the simplest familiar case by employing the above rather unfamiliar method. Namely we consider the case of $z$-independent vacuum expectation value, $v(z)=v_{0}\approx 246\>{\rm GeV}$ and suppose that the $Z$-boson mass is just a number $M_{Z}(z)=m_{Z}\approx 91.2\>{\rm GeV}$. The solutions satisfying (27) and (28) are
$\phi_{1}(z;\lambda)={\rm cos}\left(z\sqrt{\lambda-m_{Z}^{\>2}}\right),\qquad\phi_{2}(z;\lambda)=\frac{1}{\sqrt{\lambda-m_{Z}^{\>2}}}{\rm sin}\left(z\sqrt{\lambda-m_{Z}^{\>2}}\right)$ (46)
for $\lambda>m_{Z}^{\>2}$ and
$\phi_{1}(z;\lambda)={\rm cosh}\left(z\sqrt{m_{Z}^{\>2}-\lambda}\right),\qquad\phi_{2}(z;\lambda)=\frac{1}{\sqrt{m_{Z}^{\>2}-\lambda}}{\rm sinh}\left(z\sqrt{m_{Z}^{\>2}-\lambda}\right)$ (47)
for $m_{Z}^{\>2}>\lambda$. Commanding the method developed in Refs. titchmarsh and kodaira1 , we are able to get
[…]
for $\lambda>m_{Z}^{\>2}$ and $d\rho_{ij}(\lambda)/d\lambda=0$ for $m_{Z}^{\>2}>\lambda$. Thus the integration region over $\lambda$ in the completeness relation (29) is automatically restricted to $\lambda>m_{Z}^{\>2}$ and we are able to get a familiar formula
$\int\sum_{i,j=1,2}\phi_{i}(z;\lambda)d\rho_{ij}(\lambda)\phi_{j}(z^{\prime};\lambda)$ $=$ $\frac{1}{2\pi}\int_{m_{Z}^{\>2}}^{+\infty}\frac{d\lambda}{\sqrt{\lambda-m_{Z}^{\>2}}}{\rm cos}\left((z-z^{\;\prime})\sqrt{\lambda-m_{Z}^{\>2}}\right)$ (51)
$=$ $\frac{1}{\pi}\int_{0}^{+\infty}dq\>{\rm cos}\left((z-z^{\;\prime})\>q\right)$
$=$ $\delta(z-z^{\>\prime})\>.$
Note that $q\equiv\sqrt{\lambda-m_{Z}^{\>2}}\geq 0$ in (51) plays the role of the $z$-component of the wave vector. This example shows clearly that our expansion by using $\phi_{i}(z;\lambda)\;(i=1,2)$ for the general $z$-dependent $M_{Z}(z)$ case is a natural generalization of the conventional Fourier expansion method.
The mode expansion (30) of the scalar field $\varphi(t,\vec{x}_{\perp},z)$ can also be cast into the familiar form
$\varphi(t,\vec{x}_{\perp},z)$ $=$ $\int_{0}^{+\infty}dq\>\int\frac{d^{2}\vec{p}_{\perp}}{\sqrt{(2\pi)^{3}2E}}\bigg{\{}$ (52)
$\qquad a_{-}(\vec{p}_{\perp},q)e^{i\vec{p}_{\perp}\cdot\vec{x}_{\perp}+iqz-iEt}+a_{+}(\vec{p}_{\perp},q)e^{i\vec{p}_{\perp}\cdot\vec{x}_{\perp}-iqz-iEt}$
$\qquad+a_{-}^{{\dagger}}(\vec{p}_{\perp},q)e^{-i\vec{p}_{\perp}\cdot\vec{x}_{\perp}-iqz+iEt}+a_{+}^{{\dagger}}(\vec{p}_{\perp},q)e^{-i\vec{p}_{\perp}\cdot\vec{x}_{\perp}+iqz+iEt}\bigg{\}}\>,$
if we define linear combinations of the operators
$a_{\pm}(\vec{p}_{\perp},q)=\sqrt{2\pi}\left\{\sqrt{\lambda-m_{Z}^{\>2}}\>\alpha_{1}(\vec{p},\lambda)\pm i\alpha_{2}(\vec{p}_{\perp},\lambda)\right\},\qquad\left(q=\sqrt{\lambda-m_{Z}^{2}}\right)\>.$ (53)
Note that $E$ in (52) is given by $E=\sqrt{|\vec{p}_{\perp}|^{2}+\lambda}=\sqrt{|\vec{p}_{\perp}|^{2}+q^{2}+m_{Z}^{\>2}}$  . The non-vanishing commutation relations of the operators (53) are confirmed to be
$\left[a_{+}(\vec{p}_{\perp},q),a_{+}^{{\dagger}}(\vec{p}_{\perp}^{\>\prime},q^{\>\prime})\right]=\delta^{2}(\vec{p}_{\perp}-\vec{p}_{\perp}^{\>\prime})\delta(q-q^{\>\prime})\>,$ (54)
$\left[a_{-}(\vec{p}_{\perp},q),a_{-}^{{\dagger}}(\vec{p}_{\perp}^{\>\prime},q^{\>\prime})\right]=\delta^{2}(\vec{p}_{\perp}-\vec{p}_{\perp}^{\>\prime})\delta(q-q^{\>\prime})\>.$ (55)
These are the ordinary relations between creation and annihilation operators, and we are led to interpret $a_{+}(\vec{p}_{\perp},q)$ and $a_{-}(\vec{p}_{\perp},q)$ as annihilation operators of a particle with momentum $(E,\vec{p}_{\perp},-q)$ and $(E,\vec{p}_{\perp},+q)$, respectively. Since we use the real functions $\phi_{i}(z;\lambda),(i=1,2)$ instead of complex ones as the basis of the mode expansion, we have necessarily to introduce two kinds of creation and annihilation operators as in (54) and (55). This situation continues to be the case in the vector field mode expansion that will be discussed later.
Incidentally the function (LABEL:eq:deltafunction) can be shown, for $M_{Z}(z)=m_{Z}$, to reduce to the familiar invariant $\Delta$-function in the following manner:
$\thinspace\Delta(t-t^{\>\prime},\vec{x}_{\perp}-\vec{x}_{\perp}^{\>\prime},z,z^{\>\prime};m_{Z})$ (56)
$=$ $-i\int\frac{d^{2}\vec{p}_{\perp}}{(2\pi)^{2}}\>e^{i\vec{p}_{\perp}\cdot(\vec{x}_{\perp}-\vec{x}^{\;\prime}_{\perp})}\frac{1}{\pi}\int_{0}^{+\infty}dq\>{\rm cos}\left((z-z^{\>\prime})q\right)\frac{1}{2E}\left\{e^{-iE(t-t^{\>\prime})}-e^{iE(t-t^{\>\prime})}\right\}$
$=$ $-i\int_{-\infty}^{+\infty}dq\>\int\frac{d^{2}\vec{p}_{\perp}}{(2\pi)^{3}2E}\>e^{i\vec{p}_{\perp}\cdot(\vec{x}_{\perp}-\vec{x}^{\;\prime}_{\perp})+iq(z-z^{\>\prime})}\left\{e^{-iE(t-t^{\>\prime})}-e^{iE(t-t^{\>\prime})}\right\}\>.$
In the last step above, the integration region $0<q<+\infty$ is extended to $\infty<q<+\infty$, owing to the symmetry property of the integrand.
4 The BRST analyses in the electroweak theory
4.1 The gauge fixing and auxiliary fields
In order to quantize the gauge fields, we have to fix the gauge, thereby introducing the Faddeev-Popov ghost fields. We now summarize the gauge-fixing procedure below, and it should be warned in advance that there is nothing new in the content of Section 4. This section is only for preparing various quantities to be used later. Lest the equations below should become too much cluttered, we define the following notations of gauge-fixing functions,
$F^{a}$ $\equiv$ $\partial^{\mu}A_{\mu}^{a}+\frac{1}{2}gv\chi^{a},\qquad(a=1,2,3)\>,$ (57)
$F^{0}$ $\equiv$ $\partial^{\mu}B_{\mu}-\frac{1}{2}g^{\>\prime}v\chi^{3}\>.$ (58)
The gauge fields associated with $SU(2)_{L}$ and $U(1)_{Y}$ gauge groups are denoted by $A_{\mu}^{a}$ $(a=1,2,3)$ and $B_{\mu}$, respectively. The gauge coupling constants are $g$ and $g^{\>\prime}$ as introduced before.
With the use of (57) and (58), the $R_{\>\xi}$ gauge with $\xi=1$ is chosen by adding the following gauge fixing terms to the original symmetric Lagrangian
${\cal L}_{\rm GF}$ $=$ $-\sum_{a=1}^{3}{\cal B}^{a}F^{a}-{\cal B}^{0}F^{0}+\frac{1}{2}\sum_{a=1}^{3}\left({\cal B}^{a}\right)^{2}+\frac{1}{2}\left({\cal B}^{0}\right)^{2}{\color[rgb]{0,0,0}{+({\rm surface\>\>terms})\>.}}$ (59)
Note that ${\cal B}^{a}$ $(a=1,2,3)$ and ${\cal B}^{0}$ are auxiliary fields of the same type as those made use of extensively in quantum electrodynamics by Nakanishi nakanishinoboru3 and by Lautrup lautrup . The surface terms in (59) are given by
$({\rm surface\>\>terms})\equiv\partial^{\mu}\left(\sum_{a=1}^{3}{\cal B}^{a}A_{\mu}^{a}+{\cal B}^{0}B_{\mu}\right)\>,$ (60)
[…]
$-\sum_{a=1}^{3}{\cal B}^{a}(\partial^{\mu}A_{\mu}^{a})-{\cal B}^{0}(\partial^{\mu}B_{\mu})+({\rm surface\>\>terms})=\sum_{a=1}^{3}(\partial^{\mu}{\cal B}^{a})A_{\mu}^{a}+(\partial^{\mu}{\cal B}^{0})B_{\mu}\>.$ (61)
Note that the momentum variables canonically conjugate to ${\cal B}^{a}$ and ${\cal B}^{0}$ turn out to be $A_{0}^{a}$ and $B_{0}$, respectively. The reason for adding these surface terms is that this makes the Lagrangian density invariant under the BRST transformation as argued in kugoojima2 . See Eq. (81) to be given later for more details.
As an additional remark, it should be mentioned that the gauge fixing Lagrangian (59) can be rewritten identically as
${\cal L}_{\rm GF}$ $\equiv$ ${\color[rgb]{0,0,0}{-}}\left(F^{+}{\cal B}^{-}+F^{-}{\cal B}^{+}\right){\color[rgb]{0,0,0}{-}}F^{Z}{\cal B}^{Z}{\color[rgb]{0,0,0}{-}}F^{A}{\cal B}^{A}+{\cal B}^{+}{\cal B}^{-}+\frac{1}{2}({\cal B}^{Z})^{2}+\frac{1}{2}({\cal B}^{A})^{2}$ (62)
${\color[rgb]{0,0,0}{+({\rm surface\>\>terms})\>,}}$
where we have introduced the following linear combinations of (57) and (58),
$F^{\pm}$ $\equiv$ $\frac{F^{1}\mp iF^{2}}{\sqrt{2}}=\partial^{\mu}W_{\mu}^{\pm}+M_{W}\>\chi^{\pm}\>,$ (63)
[…]
$F^{A}$ $\equiv$ $\frac{g^{\>\prime}F^{3}+gF^{0}}{\sqrt{g^{2}+g^{\>\prime\>2}}}=\partial^{\mu}A_{\mu}\>.$ (65)
together with the redefined auxiliary fields,
${\cal B}^{\pm}\equiv\frac{{\cal B}^{1}\mp i{\cal B}^{2}}{\sqrt{2}}\>,\qquad{\cal B}^{Z}\equiv\>\>\>\frac{g{\cal B}^{3}-g^{\>\prime}{\cal B}^{0}}{\sqrt{g^{2}+g^{\>\prime\>2}}}\>,\qquad{\cal B}^{A}\equiv\frac{g^{\>\prime}{\cal B}^{3}+g{\cal B}^{0}}{\sqrt{g^{2}+g^{\>\prime\>2}}}\>.$ (66)
We also defined $\chi^{\pm}$ in (63) by
$\chi^{\pm}=\frac{\chi^{1}\mp i\chi^{2}}{\sqrt{2}}\>.$ (67)
The point of rewriting ${\cal L}_{\rm GF}$ as in (62) is that the gauge fields in (63), (64) and (65) are arranged so that they turn out to be $W^{\pm}$-, $Z$- and the electromagnetic ($A_{\mu}$) fields which are related to the original gauge fields $A_{\mu}^{a}$ $(a=1,2,3)$ and $B_{\mu}$ via
$W_{\mu}^{\pm}\equiv\frac{A_{\mu}^{1}\mp iA_{\mu}^{2}}{\sqrt{2}}\>,\>\>\>Z_{\mu}\equiv\frac{gA_{\mu}^{3}-g^{\>\prime}B_{\mu}}{\sqrt{g^{2}+g^{\>\prime\>2}}}\>,\>\>\>A_{\mu}\equiv\frac{g^{\>\prime}A_{\mu}^{3}+gB_{\mu}}{\sqrt{g^{2}+g^{\>\prime\>2}}}\>,$ (68)
[…]
The surface terms in (62) are given by
$({\rm surface\>\>terms})=\partial^{\mu}\left(W_{\mu}^{+}{\cal B}^{-}+W_{\mu}^{-}{\cal B}^{+}+Z_{\mu}{\cal B}^{Z}+A_{\mu}{\cal B}^{A}\right)\>,$ (69)
[…]
4.2 The Faddeev-Popov ghost Lagrangian
We just follow the standard procedure to write down the Faddeev-Popov Lagrangian ${\cal L}_{FP}$ associated with the gauge fixing (59), that is,
${\cal L}_{\rm FP}$ $=$ $\sum_{a=1}^{3}(\partial^{\mu}{\overline{c}}^{a})\>({\cal D}_{\mu}c)^{a}+(\partial^{\mu}{\overline{c}}^{0})(\partial_{\mu}c^{0})+\frac{1}{4}g^{2}v\sum_{a,b,c=1}^{3}\varepsilon^{abc}\>{\overline{c}}^{a}\chi^{b}c^{c}$ (70)
$-\frac{1}{4}g^{2}v\left({\overline{c}}^{1}c^{1}+{\overline{c}}^{2}c^{2}\right)\left(v+H\right)-\frac{1}{4}v\left(g{\overline{c}}^{3}-g^{\>\prime}{\overline{c}}^{0}\right)\left(gc^{3}-g^{\>\prime}c^{0}\right)(v+H)$
$+\frac{1}{4}gg^{\>\prime}v\left\{\left({\overline{c}}^{1}c^{0}+{\overline{c}}^{0}c^{1}\right)\chi^{2}-\left({\overline{c}}^{2}c^{0}+{\overline{c}}^{0}c^{2}\right)\chi^{1}\right\}\>.$
Here the gauge covariant derivative ${\cal D}_{\mu}$ is given as usual by
$({\cal D}_{\mu}c)^{a}\equiv\partial_{\mu}c^{a}+g\varepsilon^{abc}c^{b}A_{\mu}^{c}\>.$ (71)
On the ground of argument in Refs.kugoojima1 - kugoojima5 as regard to the (anti-)hermiticity properties of ghost and anti-ghost fields, we take
$c^{a\>\dagger}=c^{a},\qquad\overline{c}^{a\>\dagger}=-\overline{c}^{a},\qquad(a=1,2,3)$
$c^{0\>\dagger}=c^{0},\qquad\overline{c}^{0\>\dagger}=-\overline{c}^{0}\>,$ (72)
for granted. (In some of literatures, $i\>\overline{c}^{a}$ and $i\>\overline{c}^{0}$ are sometimes called anti-ghost fields, and with such naming the anti-ghost fields would become hermitian.) Thanks to the properties (72), the BRST charge $Q_{B}$ to be introduced later turns out to be a hermitian operator.
If we use the expression (62) as the gauge fixing Lagrangian, we have to rewrite the gauge fields $A_{\mu}^{\>a}$ in (70) in terms of (68). We also had better make use of the redefined ghost and anti-ghost fields
$c^{(\pm)}=\frac{c^{1}\mp ic^{2}}{\sqrt{2}}\>,\hskip 28.45274ptc^{A}=\frac{g^{\>\prime}c^{3}+gc^{0}}{\sqrt{g^{2}+g^{\>\prime\>2}}}\>,\hskip 28.45274ptc^{Z}=\frac{gc^{3}-g^{\>\prime}c^{0}}{\sqrt{g^{2}+g^{\prime\>2}}}\>,$ (73)
${\overline{c}}^{(\pm)}=\frac{{\overline{c}}^{1}\mp i{\overline{c}}^{2}}{\sqrt{2}}\>,\hskip 28.45274pt{\overline{c}}^{A}=\frac{g^{\>\prime}{\overline{c}}^{3}+g{\overline{c}}^{0}}{\sqrt{g^{2}+g^{\>\prime\>2}}}\>,\hskip 28.45274pt{\overline{c}}^{Z}=\frac{g{\overline{c}}^{3}-g^{\>\prime}{\overline{c}}^{0}}{\sqrt{g^{2}+g^{\>\prime\>2}}}\>.$ (74)
We note here an identity
${\overline{c}}^{(+)}F^{-}+{\overline{c}}^{(-)}F^{+}+{\overline{c}}^{Z}F^{Z}+{\overline{c}}^{A}F^{A}=\sum_{a=1}^{3}{\overline{c}}^{a}F^{a}+{\overline{c}}^{0}F^{0}\>,$ (75)
which determines the relations (74) among the anti-ghost fields. It is almost straightforward to rewrite (70) in terms of (73) and (74) and we relegate the rewritten form of (70) in Appendix A. See also Refs. aoki1 - aoki3 .
4.3 The BRST transformation
The BRST transformation (to be denoted by $\delta_{B}$) of the gauge and Higgs doublet fields is obtained simply by replacing the $SU(2)_{L}$ and $U(1)_{Y}$ gauge transformation parameters by the ghost fields $c^{a},\>\>(a=1,2,3)$ and $c^{0}$, respectively, i.e.,
$\delta_{B}A_{\mu}^{a}=\delta\lambda\>({\cal D}_{\mu}c)^{a},\hskip 28.45274pt\delta_{B}B_{\mu}=\delta\lambda\>\partial_{\mu}c^{0},$ (76)
$\delta_{B}\Phi=\delta\lambda\left(-\frac{i}{2}g\sum_{a=1}^{3}\tau^{a}c^{a}-\frac{i}{2}g^{\>\prime}c^{0}\right)\Phi\>,$ (77)
where $\delta\lambda$ is a Grassmann parameter. The transformation rules for the ghost fields are fixed by requiring the nilpotency, i.e., $\delta_{B}^{2}A_{\mu}^{a}=0$, $\delta_{B}^{2}B_{\mu}=0$ and $\delta_{B}^{2}\Phi=0$ as follows,
$\delta_{B}c^{a}=\frac{1}{2}g\>\delta\lambda\>\varepsilon^{abc}c^{b}c^{c},\>\>\>(a=1,2,3),\hskip 22.76228pt\delta_{B}c^{0}=0\>.$ (78)
The anti-ghost fields, on the other hand, are transformed into the auxiliary fields
$\delta_{B}\overline{c}^{a}=-\delta\lambda\>{\cal B}^{a},\>\>\>(a=1,2,3)\>,\hskip 56.9055pt\delta_{B}\overline{c}^{0}=-\delta\lambda\>{\cal B}^{0}\>,$ (79)
and the properties of nilpotency, $\delta_{B}^{2}\overline{c}^{a}=0$ and $\delta_{B}^{2}\overline{c}^{0}=0$, are guaranteed by setting the following rules,
$\delta_{B}{\cal B}^{a}=0,\>\>\>(a=1,2,3)\>,\hskip 85.35826pt\delta_{B}{\cal B}^{0}=0.$ (80)
By construction, the sum of gauge fixing Lagrangian (59) and the Faddeev-Popov Lagrangian (70) is expressed as
$\delta\lambda\left({\cal L}_{\rm GF}+{\cal L}_{\rm FP}\right)$ $=$ $\delta_{B}\bigg{\{}\sum_{a=1}^{3}\overline{c}^{a}\bigg{(}F^{a}-\frac{1}{2}{\cal B}^{a}\bigg{)}+\overline{c}^{0}\bigg{(}F^{0}-\frac{1}{2}{\cal B}^{0}\bigg{)}\bigg{\}}$ (81)
$-\delta_{B}\bigg{\{}\partial^{\mu}\bigg{(}\sum_{a=1}^{3}\overline{c}^{a}A^{a}_{\mu}+\overline{c}^{0}B_{\mu}\bigg{)}\bigg{\}}\>,$
and the invariance of the total Lagrangian density (including the surface terms) under the BRST transformation is obvious by virtue of the nilpotency .
4.4 The conserved current and the conserved charge
Now that we have the BRST invariant Lagrangian, it is straightforward to derive the conserved Noether current following the standard method, namely,
$J_{B}^{\mu}$ $=$ $\sum_{a=1}^{3}\left\{(\partial^{\mu}{\cal B}^{a})c^{a}-({\cal D}^{\mu}c)^{a}{\cal B}^{a}\right\}+(\partial^{\mu}{\cal B}^{0})c^{0}-(\partial^{\mu}c^{0}){\cal B}^{0}+\frac{1}{2}g\>\sum_{a,b,c=1}^{3}\varepsilon^{abc}(\partial^{\mu}\overline{c}^{a})c^{b}c^{c}$ (82)
$-\sum_{a=1}^{3}\partial_{\nu}\left(F^{a\>\mu\nu}c^{a}\right)-\partial_{\nu}\left\{\left(\partial^{\mu}B^{\nu}-\partial^{\nu}B^{\mu}\right)c^{0}\right\}\>.$
Here use has been made of the equations of motion of gauge fields to eliminate some part of contributions of the scalar fields. The conserved charge is the space integral of $J_{B}^{0}$ and is given by
$Q_{B}=\int d^{2}\vec{x}_{\perp}\int dz\left[\sum_{a=1}^{3}\left\{\dot{{\cal B}}^{a}c^{a}-{\cal B}^{a}({\cal D}^{0}c)^{a}\right\}+\dot{{\cal B}}^{0}c^{0}-{\cal B}^{0}\dot{c}^{0}+\frac{1}{2}g\>\sum_{a,b,c=1}^{3}\varepsilon^{abc}\dot{\overline{c}}^{a}c^{b}c^{c}\right]\>,$
where we discarded the last surface terms in (82). As the property of the conserved charge we have the commutation relations
$i\left[\delta\lambda Q_{B},{\cal O}(t,\vec{x}_{\perp},z)\right]=\delta_{B}{\cal O}(t,\vec{x}_{\perp},z)\>,$ (84)
where ${\cal O}(t,\vec{x}_{\perp},z)$ denotes generic Heisenberg fields in the electroweak theory such as gauge, scalar and auxiliary fields, together with ghost and anti-ghost fields. In terms of the physical field variables, i.e., (68), (73) and (74), $\delta_{B}{\cal O}(t,\vec{x}_{\perp},z)$’s are all given in Appendix B.
We also have another conservation law of the ghost number, assigning +1+1 for the ghost fields and 1-1 for the anti-ghost fields. The conserved Noether current turns out to be
$J_{c}^{\mu}=\sum_{a=1}^{3}\left\{\overline{c}^{a}({\cal D}^{\mu}c)^{a}-\partial^{\mu}\overline{c}^{a}\>c^{a}\right\}+\overline{c}^{0}(\partial^{\mu}c^{0})-(\partial^{\mu}\overline{c}^{0})c^{0}$ (85)
and the conserved charge is
$Q_{c}=\int d\vec{x}_{\perp}dz\>J_{c}^{0}\>.$ (86)
The Kugo-Ojima’s subsidiary condition of the physical states ($|{\rm phys}\rangle$) is expressed, in terms of $Q_{B}$ and $Q_{c}$, by
$Q_{B}|{\rm phys}\rangle=0,\hskip 28.45274ptQ_{c}|{\rm phys}\rangle=0.$ (87)
Note that the fields appearing in $Q_{B}$ and $Q_{c}$ are all interacting Heisenberg fields and the commutation relations of $Q_{B}$ with asymptotic fields have to be discussed later.
5 Field quantization in the presence of the electroweak bubble wall
5.1 Asymptotic gauge fields
Let us now write down the equations of motion of asymptotic fields which are linear with respect to fields. The vector fields, $W^{\pm}_{\mu}$, $Z_{\mu}$ and $A_{\mu}$, satisfy
$\left(\square+M_{Z}^{\>2}\right)Z_{\mu}+2M_{Z}\frac{\partial_{\mu}v}{v}\chi^{3}$ $=$ $0\>,$ (88)
$\left(\square+M_{W}^{\>2}\right)W^{\pm}_{\mu}+2M_{W}\frac{\partial_{\mu}v}{v}\chi^{\pm}$ $=$ $0\>,$ (89)
$\square A_{\mu}$ $=$ $0\>.$ (90)
and we notice that unphysical scalar fields appear in (88) and (89) due to $\partial_{\mu}v\neq 0$ . Later we will assume that the vacuum expectation value $v$ depends on only one space coordinate, $z$ (i.e., $v=v(z)$), but for now we set $\partial_{\mu}v\neq 0$ $(\mu=0,1,2,3)$, keeping generality as much as possible. The masses of the gauge bosons are given as before,
$M_{Z}^{2}$ $=$ $\frac{1}{4}\left(g^{2}+g^{\>\prime\>2}\right)v^{2}\>,\hskip 28.45274ptM_{W}^{2}=\frac{1}{4}g^{2}v^{2}\>.$ (91)
and are supposed to depend on the space-time coordinates through $v$.
The linearized equations of motion of the unphysical scalar fields also contain coupling terms with the gauge field through $\partial_{\mu}v\neq 0$,
$\left(\square+M_{1}^{\>2}\right)\chi^{3}-2M_{Z}\frac{\partial^{\mu}v}{v}Z_{\mu}$ $=$ $0\>,$ (92)
$\left(\square+M_{2}^{\>2}\right)\chi^{\pm}-2M_{W}\frac{\partial^{\mu}v}{v}W_{\mu}^{\pm}$ $=$ $0\>.$ (93)
In contrast with the usual case, the mass terms $M_{1}^{2}$ and $M_{2}^{2}$ in (92) and (93) differ from the gauge bosom masses in the following way,
$M_{1}^{\>2}=M_{Z}^{\>2}+V^{\>\prime}\left(\frac{v^{2}}{2}\right)\>,\hskip 28.45274ptM_{2}^{\>2}=M_{W}^{\>2}+V^{\>\prime}\left(\frac{v^{2}}{2}\right)\>.$ (94)
We must not fail to include the second term $V^{\>\prime}\left(v^{2}/2\right)$ which will play non-trivial roles when the absence of tadpoles is considered.
Let us turn to the auxiliary fields, which are expressed in terms of the gauge and unphysical scalar fields via equations of motion,
${\cal B}^{Z}=\partial_{\mu}Z^{\mu}+M_{Z}\chi^{3}\>,\qquad{\cal B}^{\pm}=\partial_{\mu}W^{\pm\>\mu}+M_{W}\chi^{\pm}\>,\qquad{\cal B}^{A}=\partial_{\mu}A^{\mu}\>.$ (95)
[…]
$\left(\square+M_{Z}^{\>2}\right){\cal B}^{Z}+M_{Z}\left\{\frac{1}{v}\square\>v+V^{\>\prime}\left(\frac{v^{2}}{2}\right)\right\}\chi^{3}$ $=$ $0\>,$ (96)
$\left(\square+M_{W}^{\>2}\right){\cal B}^{\pm}+M_{W}\left\{\frac{1}{v}\square\>v+V^{\>\prime}\left(\frac{v^{2}}{2}\right)\right\}\chi^{\pm}$ $=$ $0\>,$ (97)
$\square\>{\cal B}^{A}$ $=$ $0\>.$ (98)
It is extremely curious to remark that once we impose the vanishing tadpole condition (11), Eqs. (96) and (97) become simple Klein-Gordon type equations,
$\left(\square+M_{Z}^{\>2}\right){\cal B}^{Z}$ $=$ $0\>,\hskip 28.45274pt\left(\square+M_{W}^{\>2}\right){\cal B}^{\pm}=0\>.$ (99)
This implies that, under the condition (11), the mode expansion of the auxiliary fields should go along the line in Section 3.
5.2 The mode-expansion and field quantization for $v=v(z)$
Let us scrutinize the $Z$-boson field’s propagation in the similar manner to the scalar field case given in Section 3.2. The $W^{\pm}$-boson field’s propagation will be treated exactly in the same way as the $Z$-boson field and its exposition need not be repeated. The equation of the electromagnetic field is unchanged as we see in (90) and will not be discussed in the rest of the present paper. Here and hereafter we assume that the vacuum expectation value depends only on the $z$-coordinate, namely,
$v=v(z)\>.$ (100)
With this setting, the wave equations of $Z^{\mu}$-field are handled in two distinct ways, one for $\mu=0,1,2$ case and the other for $\mu=3$ case.
5.2.1 The case of $Z^{\mu}$ $(\mu=0,1,2)$ fields
The equations of $Z^{\mu}$ $(\mu=0,1,2)$ fields, as given in (88),
$\left\{\square+M_{Z}(z)^{2}\right\}Z^{\>\mu}(t,\vec{x}_{\perp},z)$ $=$ $0\>,\qquad(\mu=0,1,2)$ (101)
are separated from the unphysical scalar fields and look much simpler than the $Z^{3}(t,\vec{x}_{\perp},z)$ case. The mode expansion goes exactly in the same way as in Section 3.3. Before moving to the mode expansion, however, we have to prepare a set of vectors describing the polarization of the propagating waves. Since we are dealing with three components in (101), we introduce only three basis vectors, which we choose as follows,
[…]
Note that $\varepsilon^{(0)\>\mu}$ is time-like, while $\varepsilon^{(1)\>\mu}$ and $\varepsilon^{(2)\>\mu}$ are space-like. They are mutually orthogonal and are normalized appropriately, i.e.,
$\varepsilon^{(0)}\cdot\varepsilon^{(0)}=1,\qquad\varepsilon^{(1)}\cdot\varepsilon^{(1)}=\varepsilon^{(2)}\cdot\varepsilon^{(2)}=-1\>,$ (125)
$\varepsilon^{(0)}\cdot\varepsilon^{(1)}=\varepsilon^{(1)}\cdot\varepsilon^{(2)}=\varepsilon^{(2)}\cdot\varepsilon^{(0)}=0\>.$ (126)
It should also be useful to note that these three vectors satisfy the summation formula
[…]
a=02ε(a)με(a)νε(a)ε(a)=a=02b=02ε(a)μηabε(b)ν=( 1000 (131)
where
$\eta^{ab}=\left(\begin{tabular}[]{ccc}$1$&$0$&$0$\\ $0$&$-1$&$0$\\ $0$&$0$&$-1$\end{tabular}\right)^{ab}\>,\hskip 28.45274pt(a,b=0,1,2)\>.$ (135)
Both of $\varepsilon^{(1)\>\mu}$ and $\varepsilon^{(2)\>\mu}$ are orthogonal in the four-dimensional sense to the vector $p^{\mu}\equiv(E,p_{x},p_{y},p_{z})$, i.e., $p\cdot\varepsilon^{(1)}=0$, $p\cdot\varepsilon^{(2)}=0$, whatsoever the $z$-component $p_{z}$ would be. In the three-dimensional sense, on the other hand, only $\varepsilon^{(1)\>i}(i=1,2,3)$ is orthogonal to $\vec{p}=(p_{x},p_{y},p_{z})$. In the present study of the $z$-dependent mass, however, the $z$-component $p_{z}$ of the gauge field wave vector is not well-defined and we have to refrain from using the terminology such as longitudinal or transversal polarizations. The vectors in (124) are simply mutually-orthogonal basis vectors to be used for the expansion of $Z^{\mu}\>\>\>(\mu=0,1,2)$ according to the propagation modes. The expansion is thus given by
$Z^{\>\mu}(t,\vec{x}_{\perp},z)$ $=$ $\int d\lambda\sum_{i=1,2}\int\frac{d^{2}\vec{p}_{\perp}}{\sqrt{(2\pi)^{2}\>2E}}\sum_{a=0}^{2}\bigg{\{}\alpha_{i}^{\>(a)}(\vec{p}_{\perp},\lambda)\>e^{i\vec{p}_{\perp}\cdot\vec{x}_{\perp}-iEt}$ (136)
$\thinspace+\alpha_{i}^{\>(a)\>{\dagger}}(\vec{p}_{\perp},\lambda)\>e^{-i\vec{p}_{\perp}\cdot\vec{x}_{\perp}+iEt}\bigg{\}}\phi_{i}(z;\lambda)\>\varepsilon^{(a)\>\mu}(E,\vec{p}_{\perp})\>,\quad(\mu=0,1,2)\>.$
The coefficients $\alpha_{i}^{\>(a)}(\vec{p}_{\perp},\lambda)$ and $\alpha_{i}^{\>(a)\>{\dagger}}(\vec{p}_{\perp},\lambda)$ are operators, whose commutation relations we are about to discuss.
5.2.2 Quantization of the $Z^{\mu}$ ($\mu=0,1,2$)-fields
In analogy with the scalar case, we postulate the commutation relation
$\left[\alpha^{(a)}_{i}(\vec{p}_{\perp},\lambda),\;\alpha^{(b)\>{\dagger}}_{j}(\vec{p}_{\perp}^{\>\prime},\lambda^{\>\prime})\right]$ $=$ $-\eta^{ab}\>\delta^{2}(\vec{p}_{\perp}-\vec{p}_{\perp}^{\>\prime})\>\delta(\lambda-\lambda^{\>\prime})\>\frac{d\rho_{ij}(\lambda)}{d\lambda}$
$(a,b=0,1,2;\>\>\>i,j=1,2)\>,$
where $\eta^{ab}$ was defined by (135). Although the four-dimensional Lorentz covariance is not preserved due to the bubble wall, we still have the $O(1,2)$ symmetry in the $(t,x,y)$ subspace, and we have to put $\eta^{ab}$ on the right-hand side of (5.2.2). In connection with this, let us recall that $d\rho_{ij}(\lambda)/d\lambda$ is a positive semi-definite matrix and note the sign difference on the right hand side, $\eta^{00}=-1$ in comparison with $\eta^{11}=-\eta^{22}=+1$. The states in the Hilbert space constructed by applying the creation operator $\alpha_{i}^{(0)\>{\dagger}}(\vec{p}_{\perp},\lambda)$ on the vacuum become necessarily negative norm states. We will discuss later the decoupling of negative norm states from the physical space. Here we just confirm the following formula,
$\thinspace\sum_{a=0}^{2}\sum_{b=0}^{2}\left[\alpha^{(a)}_{i}(\vec{p}_{\perp},\lambda)\varepsilon^{(a)\>\mu}(E,\vec{p}_{\perp})\;,\alpha^{(b)\>{\dagger}}_{j}(\vec{p}_{\perp}^{\>\prime},\lambda^{\>\prime})\varepsilon^{(b)\>\nu}(E^{\>\prime},\vec{p}_{\perp}^{\>\prime})\right]$ (138)
$=$ $-\eta^{\mu\nu}\>\delta^{2}(\vec{p}_{\perp}-\vec{p}_{\perp}^{\>\prime})\>\delta(\lambda-\lambda^{\>\prime})\>\frac{d\rho_{ij}(\lambda)}{d\lambda}\>,\hskip 28.45274pt(\mu,\nu=0,1,2)\>,$
where $E^{\>\prime}=\sqrt{\lambda^{\>\prime}+|\vec{p}_{\perp}^{\>\prime}|^{2}}$ and use has been made of the summation formula (131).
Now that we are equipped with the formula (138), we are able to confirm the four-dimensional commutation relation of the $Z^{\mu}(t,\vec{x}_{\perp},z)$ field of (136) in the following form
$\left[Z^{\>\mu}(t,\vec{x}_{\perp},z),Z^{\>\nu}(t^{\>\prime},\vec{x}_{\perp}^{\>\prime},z^{\>\prime})\right]=-i\>\Delta(t-t^{\>\prime},\vec{x}_{\perp}-\vec{x}_{\perp}^{\>\prime},z,z^{\>\prime};M_{Z}(*))\>\eta^{\mu\nu}\>,$ (139)
$(\mu,\nu=0,1,2)\>.$
The equal time commutation relations are also easily deduced from above with the help of the formulas (41) - (43), i.e.,
$\left[Z^{\>\mu}(t,\vec{x}_{\perp},z),Z^{\>\nu}(t,\vec{x}_{\perp}^{\>\prime},z^{\>\prime})\right]$ $=$ $0\;,$ (140)
$\left[\dot{Z}^{\>\mu}(t,\vec{x}_{\perp},z),Z^{\>\nu}(t,\vec{x}_{\perp}^{\>\prime},z^{\>\prime})\right]$ $=$ $i\>\delta^{2}(\vec{x}_{\perp}-\vec{x}_{\perp}^{\>\prime})\delta(z-z^{\>\prime})\>\eta^{\mu\nu}\>,$ (141)
$\left[\dot{Z}^{\>\mu}(t,\vec{x}_{\perp},z),\dot{Z}^{\>\nu}(t,\vec{x}_{\perp}^{\>\prime},z^{\>\prime})\right]$ $=$ $0\;,\hskip 51.21504pt(\mu,\nu=0,1,2)\>.$
5.2.3 Coupled equations of $Z^{3}(t,\vec{x}_{\perp},z)$ and $\chi^{3}(t,\vec{x}_{\perp},z)$
Let us now turn to the equations of the $\mu=3$ component $Z^{3}(t,\vec{x}_{\perp},z)=-Z_{3}(t,\vec{x}_{\perp},z)$ and $\chi^{3}(t,\vec{x}_{\perp},z)$ which are, according to (88) and (92), coupled equations
$\left\{\square+M_{Z}(z)^{2}\right\}Z^{3}(t,\vec{x}_{\perp},z)-2M_{Z}(z)\frac{v^{\prime}(z)}{v(z)}\chi^{3}(t,\vec{x}_{\perp},z)$ $=$ $0\>,$ (142)
$\left\{\square+M_{1}(z)^{2}\right\}\chi^{3}(t,\vec{x}_{\perp},z)-2M_{Z}(z)\frac{v^{\prime}(z)}{v(z)}Z^{3}(t,\vec{x}_{\perp},z)$ $=$ $0\>,$ (143)
and therefore we have to take a different approach from the previous section. Here again we look for solutions of the plane wave type $e^{i(\vec{p}_{\perp}\cdot\vec{x}_{\perp}-Et)}$ in the $t$- and $\vec{x}_{\perp}$-directions and the coupled equations can be put in the form of the one-dimensional Schrödinger type equation with two-component wave functions,
$\left\{-\frac{\partial^{2}}{\partial z^{2}}+U(z)\right\}\left(\begin{tabular}[]{c}$Z^{3}(t,\vec{x}_{\perp},z)$\\ \\ $\chi^{3}(t,\vec{x}_{\perp},z)$\end{tabular}\right)=\lambda\>\left(\begin{tabular}[]{c}$Z^{3}(t,\vec{x}_{\perp},z)$\\ \\ $\chi^{3}(t,\vec{x}_{\perp},z)$\end{tabular}\right),\;\left(\lambda=E^{2}-|\vec{p}_{\perp}|^{2}\right).$ (150)
The "potential term" $U(z)$ is a 2 $\times$ 2 matrix whose components are given as follows,
[…]
Note that $U(z)$ is a symmetric matrix and that the differential operator on the left hand side of (150) is a self-adjoint operator.
With regard to (150), let us recall that Kodaira investigated in kodaira2 the eigenfunction expansion method associated with any even order self-adjoint differential operators as a straightforward generalization of his previous work kodaira1 . As he argued in kodaira2 , his result can be readily extended to the case of simultaneous differential equations. More specifically speaking, second-order self-adjoint differential operators for two component wave functions as given in (150) can be studied in the same way as in the case of fourth-order differential equations for a single wave function. In fact we can go in the following way.
Let us consider the following ordinary differential equation in the infinite region $\infty<z<+\infty$,
$\left\{-\frac{d^{2}}{dz^{2}}+U(z)\right\}\Phi^{(k)}(z;\lambda)=\lambda\>\Phi^{(k)}(z;\lambda)\>,\hskip 28.45274pt(k=1,\cdots,4)\;,$ (155)
where $\Phi^{(k)}(z;\lambda)$ has two components of wave functions,
$\Phi^{(k)}(z;\lambda)=\left(\begin{tabular}[]{c}$\Phi_{1}^{\>(k)}(z;\lambda)$\\ \\ $\Phi_{2}^{\>(k)}(z;\lambda)$\end{tabular}\right),\qquad(k=1,\cdots,4)\>.$ (159)
The superscript “(k)(k)" attached to $\Phi^{(k)}(z;\lambda)$ discriminates the initial conditions imposed at $z=0$ which are listed below,
[…]
Here differentiation with respect to $z$ is denoted by prime ($\prime$).
Every solution to (150) can be expressed as a linear combination of these four fundamental solutions, and the completeness is guaranteed by the formula
$\int\sum_{k,l=1}^{4}\Phi^{(k)}_{\sigma}(z;\lambda)d\widehat{\rho}_{kl}(\lambda)\Phi^{(l)}_{\tau}(z^{\>\prime};\lambda)=\delta_{\sigma\tau}\delta(z-z^{\>\prime})\>,\qquad(\sigma,\tau=1,2)\>.$ (188)
Here $\widehat{\rho}_{kl}(\lambda)$ is a 4 $\times$ 4 matrix and describes the spectrum of $\lambda$. The integration $d\widehat{\rho}_{kl}(\lambda)$ is, precisely speaking, the Stieltjes integral. We, however, assume simply that the $\lambda$-spectrum is continuous and therefore we regard (188) as the usual Riemann integration. We are thus led to the expansion of the pair of the fields in (150) in the following way,
[…]
$\hskip 28.45274pt+\beta_{k}^{\>{\dagger}}(\vec{p}_{\perp},\lambda)e^{-i\vec{p}_{\perp}\cdot\vec{x}_{\perp}+iEt}\bigg{\}}\Phi^{(k)}(z;\lambda)\>.$
The coefficients $\beta_{k}(\vec{p}_{\perp},\lambda)$ and $\beta_{k}^{\>{\dagger}}(\vec{p}_{\perp},\lambda)$ are operators, which are going to describe, respectively, the annihilation and creation of wave modes of $Z^{3}(t,\vec{x}_{\perp},z)$ and at the same time those of $\chi^{3}(t,\vec{x}_{\perp},z)$.
5.2.4 Quantization of $Z^{3}(t,\vec{x}_{\perp},z)$ and $\chi^{3}(t,\vec{x}_{\perp},z)$ fields
We now propose to set up the following commutation relations
$\left[\beta_{k}(\vec{p}_{\perp},\lambda),\beta^{\>{\dagger}}_{l}(\vec{p}_{\perp}^{\>\prime},\lambda^{\>\prime})\right]$ $=$ $\delta^{2}(\vec{p}_{\perp}-\vec{p}_{\perp}^{\>\prime})\delta(\lambda-\lambda^{\>\prime})\frac{d\widehat{\rho}_{kl}(\lambda)}{d\lambda}$ (194)
analogously to (5.2.2). Since $d\widehat{\rho}_{kl}(\lambda)/d\lambda$ is a symmetric positive definite matrix, this commutation relation does not generate negative norm states when $\beta_{k}^{{\dagger}}(\vec{p}_{\perp},\lambda)$ is regarded as a creation operator of the wave mode labeled by $(\vec{p}_{\perp},\lambda,k)$.The four-dimensional commutation relations among $Z^{3}(t,\vec{x}_{\perp},z)$ and $\chi^{3}(t,\vec{x}_{\perp},z)$ are easily worked out as follows,
$\left[Z^{3}(t,\vec{x}_{\perp},z),\>Z^{3}(t^{\>\prime},\vec{x}_{\perp}^{\>\prime},z^{\>\prime})\right]=i\widehat{\Delta}_{11}(t-t^{\>\prime},\vec{x}_{\perp}-\vec{x}_{\perp}^{\>\prime},z,z^{\>\prime};U(*))\>,$ (195)
$\left[Z^{3}(t,\vec{x}_{\perp},z),\>\chi^{3}(t^{\>\prime},\vec{x}_{\perp}^{\>\prime},z^{\>\prime})\right]=i\widehat{\Delta}_{12}(t-t^{\>\prime},\vec{x}_{\perp}-\vec{x}_{\perp}^{\>\prime},z,z^{\>\prime};U(*))\>,$ (196)
$\left[\chi^{3}(t,\vec{x}_{\perp},z),\>Z^{3}(t^{\>\prime},\vec{x}_{\perp}^{\>\prime},z^{\>\prime})\right]=i\widehat{\Delta}_{21}(t-t^{\>\prime},\vec{x}_{\perp}-\vec{x}_{\perp}^{\>\prime},z,z^{\>\prime};U(*))\;,$ (197)
$\left[\chi^{3}(t,\vec{x}_{\perp},z),\>\chi^{3}(t^{\>\prime},\vec{x}_{\perp}^{\>\prime},z^{\>\prime})\right]=i\widehat{\Delta}_{22}(t-t^{\>\prime},\vec{x}_{\perp}-\vec{x}_{\perp}^{\>\prime},z,z^{\>\prime};U(*))\;.$ (198)
Here we have newly defined a $2\times 2$ matrix function,
$\thinspace\widehat{\Delta}_{\sigma\tau}(t-t^{\>\prime},\vec{x}_{\perp}-\vec{x}_{\perp}^{\>\prime},z,z^{\>\prime};U(*))$ (199)
$=$ $-i\int\frac{d^{2}\vec{p}_{\perp}}{(2\pi)^{2}}\>e^{i\vec{p}_{\perp}\cdot(\vec{x}_{\perp}-\vec{x}^{\;\prime}_{\perp})}\sum_{k,l}\int\Phi_{\sigma}^{(k)}(z;\lambda)\frac{d\widehat{\rho}_{kl}(\lambda)}{2E}\Phi_{\tau}^{(l)}(z^{\>\prime};\lambda)$
$\quad\times\left\{e^{-iE(t-t^{\>\prime})}-e^{iE(t-t^{\>\prime})}\right\}\>,\hskip 28.45274pt(\tau,\sigma=1,2)\>.$
Apparently (199) satisfies the differential equations
$\left\{\square+U(z)\right\}\left(\begin{tabular}[]{cc}$\widehat{\Delta}_{11}$&$\widehat{\Delta}_{12}$\\ $\widehat{\Delta}_{21}$&$\widehat{\Delta}_{22}$\end{tabular}\right)=0\>,\qquad\left\{\square^{\>\prime}+U(z^{\prime})\right\}\left(\begin{tabular}[]{cc}$\widehat{\Delta}_{11}$&$\widehat{\Delta}_{21}$\\ $\widehat{\Delta}_{12}$&$\widehat{\Delta}_{22}$\end{tabular}\right)=0\;,$ (204)
together with the characteristic properties at $t=t^{\>\prime}$,
$\widehat{\Delta}_{\sigma\tau}\left(0,\vec{x}_{\perp}-\vec{x}^{\>\prime}_{\perp},z,z^{\>\prime};U(*)\right)=0\>,$ (205)
$\frac{\partial}{\partial t}\widehat{\Delta}_{\sigma\tau}\left(t-t^{\>\prime},\vec{x}_{\perp}-\vec{x}^{\>\prime}_{\perp},z,z^{\>\prime};U(*)\right)\Big{|}_{t=t^{\>\prime}}=-\delta_{\sigma\tau}\delta^{2}(\vec{x}_{\perp}-\vec{x}^{\>\prime})\delta(z-z^{\>\prime})\>,$ (206)
$\frac{\partial}{\partial t^{\>\prime}}\frac{\partial}{\partial t}\widehat{\Delta}_{\sigma\tau}\left(t-t^{\>\prime},\vec{x}_{\perp}-\vec{x}^{\>\prime}_{\perp},z,z^{\>\prime};U(*)\right)\Big{|}_{t=t^{\>\prime}}=0\>.$ (207)
The formula (206) leads us to derive the equal-time commutation relations
$\left[\dot{Z}^{3}(t,\vec{x}_{\perp},z),\>Z^{3}(t,\vec{x}_{\perp}^{\>\prime},z^{\>\prime})\right]$ $=$ $-i\delta^{2}(\vec{x}_{\perp}-\vec{x}_{\perp}^{\>\prime})\delta(z-z^{\>\prime})\;,$ (208)
$\left[\dot{Z}^{3}(t,\vec{x}_{\perp},z),\>\chi^{3}(t,\vec{x}_{\perp}^{\>\prime},z^{\>\prime})\right]$ $=$ $0\;,$ (209)
$\left[\dot{\chi}^{3}(t,\vec{x}_{\perp},z),\>Z^{3}(t,\vec{x}_{\perp}^{\>\prime},z^{\>\prime})\right]$ $=$ $0\;,$ (210)
$\left[\dot{\chi}^{3}(t,\vec{x}_{\perp},z),\>\chi^{3}(t,\vec{x}_{\perp}^{\>\prime},z^{\>\prime})\right]$ $=$ $-i\delta^{2}(\vec{x}_{\perp}-\vec{x}_{\perp}^{\>\prime})\delta(z-z^{\>\prime})\;.$ (211)
5.3 The auxiliary field ${\cal B}^{Z}(t,\vec{x}_{\perp},z)$
Towards the end of Section 4.1, it was mentioned that $Z^{0}(t,\vec{x}_{\perp},z)$ is the momentum variable canonically conjugate to the auxiliary field ${\cal B}^{Z}(t,\vec{x}_{\perp},z)$, and let us look at this point further. Expressing ${\cal B}^{Z}(t,\vec{x}_{\perp},z)$ by the formula (95) we are able to compute the commutator
[…]
$=$ $-i\delta^{2}(\vec{x}_{\perp}-\vec{x}_{\perp}^{\>\prime})\delta(z-z^{\>\prime})\>,$
which is a desirable relation. We also have to confirm that ${\cal B}^{Z}(t,\vec{x}_{\perp},z)$ is a dynamical variable independent of $\chi^{3}(t,\vec{x}_{\perp},z)$ and the commuting property
$\left[\chi^{3}(t,\vec{x}_{\perp},z),\>\>{\cal B}^{Z}(t,\vec{x}_{\perp}^{\>\prime},z^{\>\prime})\right]=0$ (213)
follows from (198) and (205). The formula (211), on the other hand, tells us the relation
$\left[\dot{\chi}^{3}(t,\vec{x}_{\perp},z),\>\>{\cal B}^{Z}(t,\vec{x}_{\perp}^{\>\prime},z^{\>\prime})\right]=-i\delta^{2}(\vec{x}_{\perp}-\vec{x}_{\perp}^{\>\prime})\delta(z-z^{\>\prime})\>M_{Z}(z)\>.$ (214)
Recall that the momentum variable canonically conjugate to $\chi^{3}(t,\vec{x}_{\perp},z)$ is not $\dot{\chi}^{3}(t,\vec{x}_{\perp},z)$, but is given by
$\Pi_{\chi^{3}}(t,\vec{x}_{\perp},z)\equiv\dot{\chi}^{3}(t,\vec{x}_{\perp},z)-M_{Z}(z)\>Z^{0}(t,\vec{x}_{\perp},z)\>.$ (215)
Thanks to the second term in (215), the commuting property
$\left[\Pi_{\chi^{3}}(t,\vec{x}_{\perp},z),\>\>{\cal B}^{Z}(t,\vec{x}_{\perp}^{\>\prime},z^{\>\prime})\right]=0$ (216)
is guaranteed.
Let us now move to the commutation relations of ${\cal B}^{Z}(t,\vec{x}_{\perp},z)$ with itself. We have already listed up all the non-vanishing equal time commutation relations involving $Z^{\mu}(t,\vec{x}_{\perp},z)$ and $\chi^{3}(t,\vec{x}_{\perp},z)$ and it is straightforward to confirm the vanishing of the following equal time commutators,
$\left[{\cal B}^{Z}(t,\vec{x}_{\perp},z),\>\>{\cal B}^{Z}(t,\vec{x}_{\perp}^{\>\prime},z^{\>\prime})\right]=0\>,\qquad\left[\dot{{\cal B}}^{Z}(t,\vec{x}_{\perp},z),\>\>{\cal B}^{Z}(t,\vec{x}_{\perp}^{\>\prime},z^{\>\prime})\right]=0\>.$ (217)
As for the four-dimensional commutator of ${\cal B}^{Z}(t,\vec{x}_{\perp},z)$, we would get a lengthy formula containing the generalized invariant $\Delta$-functions of (LABEL:eq:deltafunction) and (199). It is nonetheless still possible to prove the vanishing of the four-dimensional commutator
$\left[{\cal B}^{Z}(t,\vec{x}_{\perp},z),\>\>{\cal B}^{Z}(t^{\>\prime},\vec{x}_{\perp}^{\>\prime},z^{\>\prime})\right]=0\>,$ (218)
by using (217) and the Klein-Gordon type equation of motion (99) of ${\cal B}^{Z}(t,\vec{x}_{\perp},z)$. The vanishing property of (218) is due to the cancellation between two contributions that generate positive and negative norm states. This is similar to the circumstance in the Abelian Higgs model nakanishinoboru1 - nishijima1 . The technique of proving (218) is given in Appendix C in order to avoid digressing from the main stream of the present paper.
5.4 The ghost and anti-ghost fields
The asymptotic ghost and anti-ghost fields in the $R_{\xi}$ gauge with $\xi=1$ satisfy the Klein-Gordon type equations, as can be seen from the Faddeev-Popov Lagrangian (283) and (284) in Appendix A, expressed in terms of (73) and (74), namely,
$\left(\square+M_{W}^{2}\right)c^{(\pm)}=0,\hskip 28.45274pt\left(\square+M_{Z}^{2}\right)c^{Z}=0,\hskip 28.45274pt\square\>c^{A}=0,$ (219)
$\left(\square+M_{W}^{2}\right)\overline{c}^{(\pm)}=0,\hskip 28.45274pt\left(\square+M_{Z}^{2}\right)\overline{c}^{Z}=0,\hskip 28.45274pt\square\>\overline{c}^{A}=0\>.$ (220)
The mode expansion of these fields goes without any change as in Section 3.3. Since we focus our attention only to the $Z$-boson sector, we write down the mode expansion only of $c^{Z}(t,\vec{x}_{\perp},z)$ and $\overline{c}^{Z}(t,\vec{x}_{\perp},z)$,
$c^{Z}(t,\vec{x}_{\perp},z)$ $=$ $\int d\lambda\sum_{i=1,2}\int\frac{d^{2}\vec{p}_{\perp}}{\sqrt{(2\pi)^{2}\>2E}}$ (221)
$\times\bigg{\{}\gamma_{i}(\vec{p}_{\perp},\lambda)e^{i\vec{p}_{\perp}\cdot\vec{x}_{\perp}-iEt}+\gamma_{i}^{{\dagger}}(\vec{p}_{\perp},\lambda)e^{-i\vec{p}_{\perp}\cdot\vec{x}_{\perp}+iEt}\bigg{\}}\phi_{i}(z;\lambda)\>,$
$i\>\overline{c}^{Z}(t,\vec{x}_{\perp},z)$ $=$ $\int d\lambda\sum_{i=1,2}\int\frac{d^{2}\vec{p}_{\perp}}{\sqrt{(2\pi)^{2}\>2E}}$ (222)
$\times\bigg{\{}\overline{\gamma}_{i}(\vec{p}_{\perp},\lambda)e^{i\vec{p}_{\perp}\cdot\vec{x}_{\perp}-iEt}+\overline{\gamma}_{i}^{\>{\dagger}}(\vec{p}_{\perp},\lambda)e^{-i\vec{p}_{\perp}\cdot\vec{x}_{\perp}+iEt}\bigg{\}}\phi_{i}(z;\lambda)\>.$
Here let us recall the hermiticity of ghost and anti-hermiticity of anti-ghost fields, (72). We have therefore put i``\>i\>" on the left hand side of (222) .
The anti-commutation relations of the coefficients in (221) and (222) that we postulate here are given by
$\left\{\gamma_{i}(\vec{p}_{\perp},\lambda),\overline{\gamma}_{j}^{\>{\dagger}}(\vec{p}_{\perp}^{\>\prime},\lambda^{\>\prime})\right\}$ $=$ $i\>\delta^{2}(\vec{p}_{\perp}-\vec{p}_{\perp}^{\>\prime})\delta(\lambda-\lambda^{\>\prime})\frac{d\rho_{ij}(\lambda)}{d\lambda}\>,$
$\left\{\overline{\gamma}_{i}(\vec{p}_{\perp},\lambda),\gamma_{j}^{\>{\dagger}}(\vec{p}_{\perp}^{\>\prime},\lambda^{\>\prime})\right\}$ $=$ $-i\>\delta^{2}(\vec{p}_{\perp}-\vec{p}_{\perp}^{\>\prime})\delta(\lambda-\lambda^{\>\prime})\frac{d\rho_{ij}(\lambda)}{d\lambda}\>,\quad(i,j=1,2)\>.$ (223)
All other anti-commutators should vanish. With (223), we can compute the four-dimensional anti-commutaion relations between ghost and anti-ghost fields as
$\left\{c^{Z}(t,\vec{x}_{\perp},z),\overline{c}^{Z}(t^{\>\prime},\vec{x}_{\perp}^{\>\prime},z^{\>\prime})\right\}=i\Delta(t-t^{\>\prime},\vec{x}_{\perp}-\vec{x}_{\perp}^{\>\prime};M_{Z}(*))\>.$ (224)
Thanks to the property (42) of our $\Delta$-function, the equal-time anti-commutaion relations
$\left\{\dot{c}^{Z}(t,\vec{x}_{\perp},z),\overline{c}^{Z}(t,\vec{x}_{\perp}^{\>\prime},z^{\>\prime})\right\}$ $=$ $-i\>\delta(\vec{x}_{\perp}-\vec{x}_{\perp}^{\>\prime})\delta(z-z^{\>\prime})\>,$ (225)
$\left\{c^{Z}(t,\vec{x}_{\perp},z),\dot{\overline{c}}^{Z}(t,\vec{x}_{\perp}^{\>\prime},z^{\>\prime})\right\}$ $=$ $i\>\delta(\vec{x}_{\perp}-\vec{x}_{\perp}^{\>\prime})\delta(z-z^{\>\prime})\>,$ (226)
follow from (224) immediately.
6 More about the auxiliary field ${\cal B}^{Z}(t,\vec{x}_{\perp},z)$
In Section 5, we carried out the mode expansion of $Z^{\mu}(t,\vec{x}_{\perp},z)$ and $\chi^{3}(t,\vec{x}_{\perp},z)$ as given in (136) and (193). Since the auxiliary field ${\cal B}^{Z}(t,\vec{x}_{\perp},z)$ is expressed as a linear combination of these fields as in (95), the mode expansion of ${\cal B}^{Z}(t,\vec{x}_{\perp},z)$ is also already at our hand. Now by looking at the mode expansion of ${\cal B}^{Z}(t,\vec{x}_{\perp},z)$ carefully, we will show in Sections 6.1 and 6.2 that the creation operators contained in ${\cal B}^{Z}(t,\vec{x}_{\perp},z)$, when applied on the vacuum, give rise to only zero norm states which are unobservable. We will define another field ${\cal N}(t,\vec{x}_{\perp},z)$ in Section 6.3 and will discuss the BRST quartet mechanism in Section 6.4 by using ${\cal B}^{Z}(t,\vec{x}_{\perp},z)$ and ${\cal N}(t,\vec{x}_{\perp},z)$.
6.1 Mode expansion of ${\cal B}^{Z}(t,\vec{x}_{\perp},z)$
Let us write down the mode expansion of ${\cal B}^{Z}(t,\vec{x}_{\perp},z)$, which is, by combining (136) and (193), cast into a lengthy formula
$\thinspace{\cal B}^{Z}(t,\vec{x}_{\perp},z)$ (227)
$=$ $\partial_{\mu}Z^{\mu}(t,\vec{x}_{\perp},z)+M_{Z}(z)\chi^{3}(t,\vec{x}_{\perp},z)$
$=$ $-i\int\sqrt{\lambda}\>d\lambda\sum_{i=1,2}\int\frac{d^{2}\vec{p}_{\perp}}{\sqrt{(2\pi)^{2}2E}}$
$\qquad\times\bigg{\{}\alpha_{i}^{(0)}(\vec{p}_{\perp},\lambda)e^{i\vec{p}_{\perp}\cdot\vec{x}_{\perp}-iEt}-\alpha_{i}^{(0){\dagger}}(\vec{p}_{\perp},\lambda)e^{-i\vec{p}_{\perp}\cdot\vec{x}_{\perp}+iEt}\bigg{\}}\phi_{i}(z;\lambda)$
$+\int\>d\lambda\sum_{k=1}^{4}\int\frac{d^{2}\vec{p}_{\perp}}{\sqrt{(2\pi)^{2}2E}}\left\{\frac{d}{dz}\Phi_{1}^{(k)}(z;\lambda)+M_{Z}(z)\Phi_{2}^{(k)}(z;\lambda)\right\}$
$\qquad\times\bigg{\{}\beta_{k}(\vec{p}_{\perp},\lambda)e^{i\vec{p}_{\perp}\cdot\vec{x}_{\perp}-iEt}+\beta_{k}^{{\dagger}}(\vec{p}_{\perp},\lambda)e^{-i\vec{p}_{\perp}\cdot\vec{x}_{\perp}+iEt}\bigg{\}}\>.$
There appears an interesting and suggestive combination of $\Phi_{1}^{(k)}(z;\lambda)$ and $\Phi_{2}^{(k)}(z;\lambda)$ in the last integrand of (227), i.e.,
$\mathcal{F}^{(k)}(z;\lambda)\equiv\frac{d}{dz}\Phi_{1}^{(k)}(z;\lambda)+M_{Z}(z)\Phi_{2}^{(k)}(z;\lambda)\>,\hskip 28.45274pt(k=1,\cdots,4)\>.$ (228)
By shuffling the equation (155) for $\Phi_{1}^{(k)}(z;\lambda)$ and $\Phi_{2}^{(k)}(z;\lambda)$, we find an intriguing formula satisfied by (228),
$\left\{-\frac{d^{2}}{dz^{2}}+M_{Z}(z)^{2}-\lambda\right\}\mathcal{F}^{(k)}(z;\lambda)$ $=$ $M_{Z}(z)\left\{\frac{v^{\>\prime\prime}(z)}{v(z)}-V^{\>\prime}\left(\frac{v(z)^{2}}{2}\right)\right\}\Phi_{2}^{(k)}(z;\lambda)\>,$ (229)
$\hskip 28.45274pt(k=1,\cdots,4)\>.$
whose right hand side vanishes, if we impose the absence of tadpole, i.e., (12). In other words, under (12), the peculiar combination (228) satisfies the same equation as that in (26) and can be expressed as a linear combination of the basic solutions, $\phi_{1}(z;\lambda)$ and $\phi_{2}(z;\lambda)$. To determine the form of the linear combination, we note the values of (228) and its derivative at $z=0$ by looking at (166) - (187), , i.e.,
$\mathcal{F}^{(k)}(0;\lambda)=\mathfrak{f}_{1}^{\>(k)}\>,\hskip 56.9055pt\frac{d}{dz}\mathcal{F}^{(k)}(z;\lambda)\Bigg{|}_{z=0}=\mathfrak{f}_{2}^{\>(k)}\>,$ (230)
where
$\mathfrak{f}_{1}^{\>(1)}=0\>,\hskip 56.9055pt\mathfrak{f}_{2}^{\>(1)}=M_{Z}(0)^{2}-\lambda\>,$ (231)
$\mathfrak{f}_{1}^{\>(2)}=1\>,\hskip 56.9055pt\mathfrak{f}_{2}^{\>(2)}=0\>,$ (232)
$\mathfrak{f}_{1}^{\>(3)}=M_{Z}(0)\>,\hskip 31.2982pt\mathfrak{f}_{2}^{\>(3)}=-M_{Z}(0)\frac{v^{\>\prime}(0)}{v(0)}\>,$ (233)
$\mathfrak{f}_{1}^{\>(4)}=0\>,\hskip 56.9055pt\mathfrak{f}_{2}^{\>(4)}=M_{Z}(0)\>.$ (234)
Here use has been made of the derivative formula of (228),
$\frac{d}{dz}\mathcal{F}^{(k)}(z;\lambda)$ $=$ $\left\{M_{Z}(z)^{2}-\lambda\right\}\Phi_{1}^{(k)}(z;\lambda)-M_{Z}(z)\frac{v^{\>\prime}(z)}{v(z)}\Phi_{2}^{(k)}(z;\lambda)$ (235)
$+M_{Z}(z)\frac{d}{dz}\Phi_{2}^{(k)}(z;\lambda)\>.$
Comparing the initial value conditions at $z=0$ in (27) and (28) with the above, we are led to the following equality
$\mathcal{F}^{(k)}(z;\lambda)=\sum_{i=1}^{2}\mathfrak{f}_{i}^{(k)}\phi_{i}(z;\lambda)\>.$ (236)
Taking (236) into account, the mode expansion (227) is rendered much simpler, i.e.,
${\cal B}^{Z}(t,\vec{x}_{\perp},z)$ $=$ $\int\>d\lambda\int\frac{d^{2}\vec{p}_{\perp}}{\sqrt{(2\pi)^{2}2E}}$ (237)
$\thinspace\times\sum_{i=1,2}\bigg{\{}\mathfrak{b}_{i}(\vec{p}_{\perp},\lambda)e^{i\vec{p}_{\perp}\cdot\vec{x}_{\perp}-iEt}+\mathfrak{b}_{i}^{{\dagger}}(\vec{p}_{\perp},\lambda)e^{-i\vec{p}_{\perp}\cdot\vec{x}_{\perp}+iEt}\bigg{\}}\phi_{i}(z;\lambda)\;,$
where the coefficients are defined by
$\mathfrak{b}_{i}(\vec{p}_{\perp},\lambda)=-i\sqrt{\lambda}\>\alpha_{i}^{(0)}(\vec{p}_{\perp},\lambda)+\sum_{k=1}^{4}\>\mathfrak{f}_{i}^{(k)}\beta_{k}(\vec{p}_{\perp},\lambda)\>,\qquad(i=1,2)\>.$ (238)
The simplification of the mode expansion (237) in comparison with (227) is, however, not quite unexpected. We have already noticed before with reference to (96) and (99), that the auxiliary field ${\cal B}^{Z}(t,\vec{x}_{\perp},z)$ obeys the Klein-Gordon type equation under the condition of (11), and we have actually foreseen there that the mode expansion of ${\cal B}^{Z}(t,\vec{x}_{\perp},z)$ should go along the line in Section 3. The expansion in (237) is consistent with the previous observation.
6.2 Connection between $d\rho_{ij}(\lambda)$ and $d\widehat{\rho}_{kl}(\lambda)$
Looking at the two terms on the right hand side of (238), we notice immediately that, while $\beta_{k}^{{\dagger}}(\vec{p}_{\perp},\lambda)$ generates positive norm states when acting on the vacuum, the other $\alpha_{i}^{(0)\>{\dagger}}(\vec{p}_{\perp},\lambda)$ introduces negative norm states on the contrary. Such a mixture of qualitatively distinct type of operators is characteristic in the definition of $\mathfrak{b}_{i}(\vec{p}_{\perp},\lambda)$ and it it extremely interesting to compute its commutator with $\mathfrak{b}^{{\dagger}}(\vec{p}_{\perp}^{\>\prime},\lambda^{\>\prime})$. On applying the formulae (5.2.2) and (194), we get the following provisional result
$\thinspace\left[\mathfrak{b}_{i}(\vec{p}_{\perp},\lambda),\mathfrak{b}^{{\dagger}}_{j}(\vec{p}_{\perp}^{\>\prime},\lambda^{\>\prime})\right]$ (239)
$=$ $\sqrt{\lambda}\sqrt{\lambda^{\>\prime}}\left[\alpha_{i}^{(0)}(\vec{p}_{\perp},\lambda),\alpha_{j}^{(0)\>{\dagger}}(\vec{p}_{\perp}^{\>\prime},\lambda^{\>\prime}),\right]+\sum_{k,l=1}^{4}\mathfrak{f}_{i}^{(k)}\mathfrak{f}_{j}^{(l)}\left[\beta_{k}(\vec{p}_{\perp},\lambda),\beta_{l}^{{\dagger}}(\vec{p}_{\perp}^{\>\prime},\lambda^{\>\prime})\right]$
$=$ $\left\{-\lambda\frac{d\rho_{ij}(\lambda)}{d\lambda}+\sum_{k,l=1}^{4}\mathfrak{f}_{i}^{(k)}\frac{d\widehat{\rho}_{kl}(\lambda)}{d\lambda}\mathfrak{f}_{j}^{(l)}\right\}\delta^{2}(\vec{p}_{\perp}-\vec{p}_{\perp}^{\>\prime})\delta(\lambda-\lambda^{\>\prime})\>.$
We would be unable to go further from (239) without the knowledge on the connection between $d\rho_{ij}(\lambda)$ and $d\widehat{\rho}_{kl}(\lambda)$.
It is, however, more than likely that there must exist a close connection between $d\rho_{ij}(\lambda)$ and $d\widehat{\rho}_{kl}(\lambda)$, because the two differential equations, (26) and (150), are both given in terms of the common function $v(z)$ if we replace $V^{\>\prime}(v^{2}/2)$ in (94) by $v^{\>\prime\prime}(z)/v(z)$ by virtue of (12). The algorithm of computing $d\rho_{ij}(\lambda)$ and $d\widehat{\rho}_{kl}(\lambda)$ has in principle been established by the work of Titchmarsh titchmarsh and of Kodaira kodaira1 - kodaira2 , and it is most desirable to look at the connection along the line of their work. For now, however, we employ a short-cut method without going into too much details of mathematical complexities.
Let us now study the following integration
$\sum_{k,l=1}^{4}\int d\lambda\>{\cal F}^{(k)}(z;\lambda)\frac{d\widehat{\rho}_{kl}(\lambda)}{d\lambda}{\cal F}^{(l)}(z^{\>\prime};\lambda)$ $=$ $\sum_{k,l=1}^{4}\sum_{i,j=1}^{2}\int d\lambda\>\mathfrak{f}_{i}^{(k)}\phi_{i}(z;\lambda)\frac{d\widehat{\rho}_{kl}(\lambda)}{d\lambda}\mathfrak{f}_{j}^{(l)}\phi_{j}(z^{\>\prime};\lambda)\>.$
Putting the definition (228) of ${\cal F}^{(k)}(z;\lambda)$ into the left hand side, and using the integration formula (188), we arrive at a concise formula
$\thinspace\sum_{k,l=1}^{4}\int d\lambda\>{\cal F}^{(k)}(z;\lambda)\frac{d\widehat{\rho}_{kl}(\lambda)}{d\lambda}{\cal F}^{(l)}(z^{\>\prime};\lambda)$ (241)
$=$ $\sum_{k,l=1}^{4}\int d\lambda\>\left\{\frac{d}{dz}\Phi_{1}^{(k)}(z;\lambda)+M_{Z}(z)\Phi_{2}^{(k)}(z;\lambda)\right\}\frac{d\widehat{\rho}_{kl}(\lambda)}{d\lambda}$
$\hskip 42.67912pt\times\left\{\frac{d}{dz^{\>\prime}}\Phi_{1}^{(l)}(z^{\>\prime};\lambda)+M_{Z}(z^{\>\prime})\Phi_{2}^{(l)}(z^{\>\prime};\lambda)\right\}$
$=$ $\left\{\frac{d}{dz}\frac{d}{dz^{\>\prime}}+M_{Z}(z)M_{Z}(z^{\>\prime})\right\}\delta(z-z^{\>\prime})$
$=$ $\left\{-\frac{d^{2}}{dz^{2}}+M_{Z}(z)^{2}\right\}\delta(z-z^{\>\prime})\>.$
In the meantime we can express the right hand side of (241) in the following way, i.e.,
$\left\{-\frac{d^{2}}{dz^{2}}+M_{Z}(z)^{2}\right\}\delta(z-z^{\>\prime})$ $=$ $\left\{-\frac{d^{2}}{dz^{2}}+M_{Z}(z)^{2}\right\}\sum_{i,j=1}^{2}\int d\lambda\>\phi_{i}(z;\lambda)\frac{d\rho_{ij}(\lambda)}{d\lambda}\phi_{j}(z^{\>\prime};\lambda)$ (242)
$=$ $\sum_{i,j=1}^{2}\int d\lambda\>\phi_{i}(z;\lambda)\>\lambda\frac{d\rho_{ij}(\lambda)}{d\lambda}\phi_{j}(z^{\>\prime};\lambda)\>.$
Combining (LABEL:eq:aboveequation0), (241) and (242), we get an identity that contains both $d\rho_{ij}(\lambda)$ and $d\widehat{\rho}_{kl}(\lambda)$,
$\sum_{k,l=1}^{4}\sum_{i,j=1}^{2}\int d\lambda\>\mathfrak{f}_{i}^{(k)}\phi_{i}(z;\lambda)\frac{d\widehat{\rho}_{kl}(\lambda)}{d\lambda}\mathfrak{f}_{j}^{(l)}\phi_{j}(z^{\>\prime};\lambda)=\sum_{i,j=1}^{2}\int d\lambda\>\phi_{i}(z;\lambda)\>\lambda\frac{d\rho_{ij}(\lambda)}{d\lambda}\phi_{j}(z^{\>\prime};\lambda)\>.$
For the equality (243) to be valid for arbitrary values of $z$ and $z^{\>\prime}$, we deduce the connection between $d\rho_{ij}(\lambda)$ and $d\widehat{\rho}_{kl}(\lambda)$,
$\sum_{k,l=1}^{4}\mathfrak{f}_{i}^{(k)}\frac{d\widehat{\rho}_{kl}(\lambda)}{d\lambda}\mathfrak{f}_{j}^{(l)}$ $=$ $\>\lambda\frac{d\rho_{ij}(\lambda)}{d\lambda}\>,\qquad(i,j=1,2)\>.$ (244)
This has a deep impact on the commutator (239). Namely, we get a vanishing commutation relation
$\left[\mathfrak{b}_{i}(\vec{p}_{\perp},\lambda),\mathfrak{b}^{{\dagger}}_{j}(\vec{p}_{\perp}^{\>\prime},\lambda^{\>\prime})\right]$ $=$ $0\>,\qquad(i,j=1,2)\>.$ (245)
This vanishing commutator indicates that the operator $\mathfrak{b}^{{\dagger}}_{i}(\vec{p}_{\perp},\lambda)$, when applied on the vacuum, generates only zero-norm states. It should also be noted that (245) is consistent with the four-dimensional commutation relation (218).
6.3 Mode expansion of ${\cal N}(t,\vec{x}_{\perp},z)$
In the usual BRST analysis in the standard electroweak theory with position-independent masses aoki1 - aoki3 , the BRST quartet fields in the $Z$-boson sector consist of ${\cal B}^{Z}(t,\vec{x}_{\perp},z)$, $c^{Z}(t,\vec{x}_{\perp},z)$, $\overline{c}^{Z}(t,\vec{x}_{\perp},z)$ and the unphysical scalar field $\chi^{3}(t,\vec{x}_{\perp},z)$. For the decoupling of the quartet members from the physical S-matrix to work well, the asymptotic fields of the quartet must satisfy the same equations of motion with the common mass term. In the presence of the electroweak bubble wall of our case, ${\cal B}^{Z}(t,\vec{x}_{\perp},z)$, $c^{Z}(t,\vec{x}_{\perp},z)$ and $\overline{c}^{Z}(t,\vec{x}_{\perp},z)$ satisfy the same Klein-Gordon type equations of motion with the common mass term $M_{Z}(z)$, namely, (99), (219) and (220), but $\chi^{3}(t,\vec{x}_{\perp},z)$ does not. We have already seen that $\chi^{3}(t,\vec{x}_{\perp},z)$ satisfies the rather complicated equation (150) coupled with $Z^{3}(t,\vec{x}_{\perp},z)$. We have to seek for something else for the fourth member of the quartet.
There are two candidates for the fourth member, that is, [^2] We use the summation symbol ($\sum$) for the repeated index $\mu$, only when the summation is taken over $\mu=0,1,2$.
$\sum_{\mu=0,1,2}\partial_{\mu}Z^{\mu}(t,\vec{x}_{\perp},z),\hskip 28.45274pt\frac{\partial}{\partial z}Z^{3}(t,\vec{x}_{\perp},z)+M_{Z}(z)\chi^{3}(t,\vec{x}_{\perp},z),$ (246)
whose asymptotic fields both satisfy the Klein-Gordon type equations of motion. We can think of various linear combinations of the above two, but for definiteness we would like to take
$\thinspace{\cal N}(t,\vec{x}_{\perp},z)$ $\equiv$ $-\sum_{\mu=0,1,2}\partial_{\mu}Z^{\mu}(t,\vec{x}_{\perp},z)+\frac{\partial}{\partial z}Z^{3}(t,\vec{x}_{\perp},z)+M_{Z}(z)\chi^{3}(t,\vec{x}_{\perp},z)$ (247)
as the fourth member of the quartet. Actually in the absence of the tadpole, the asymptotic field of (247) satisfies the Klein-Gordon type equation of motion with the mass term $M_{Z}(z)$,
$\left\{\square+M_{Z}(z)^{2}\right\}{\cal N}(t,\vec{x}_{\perp},z)$ $=$ $M_{Z}(z)\left\{\frac{v^{\>\prime\prime}(z)}{v(z)}-V^{\>\prime}\left(\frac{v(z)^{2}}{2}\right)\right\}\chi^{3}(t,\vec{x}_{\perp},z)$ (248)
$=$ $0\>.$
Note that the definition (247) differs from that of ${\cal B}^{Z}(t,\vec{x}_{\perp},z)$ in (95) only by the minus sign in front of the first term in (247) and some of the formulas to be given below are in parallel with those of ${\cal B}^{Z}(t,\vec{x}_{\perp},z)$ . The principal reason for choosing (247), however, is that the combination of (247) makes the metric structure of the state vectors very convenient owing to the commutation relation (251) to be explained later. Other choices would work equally well but the analysis on the BRST quartet mechanism would become a little involved, although the conclusion is unaltered.
Since ${\cal N}(t,\vec{x}_{\perp},z)$ satisfies (248), its mode expansion goes in the same way as in Section 3.3 with the use of $\phi_{i}(z;\lambda),(i=1,2)$, namely,
${\cal N}(t,\vec{x}_{\perp},z)$ $=$ $\int d\lambda\int\frac{d^{2}\vec{p}_{\perp}}{\sqrt{(2\pi)^{2}\>2E}}$ (249)
$\thinspace\times\sum_{i=1,2}\bigg{\{}\mathfrak{n}_{i}(\vec{p}_{\perp},\lambda)\>e^{i\vec{p}_{\perp}\cdot\vec{x}_{\perp}-iEt}+\mathfrak{n}_{i}^{\>{\dagger}}(\vec{p}_{\perp},\lambda)\>e^{-i\vec{p}_{\perp}\cdot\vec{x}_{\perp}+iEt}\bigg{\}}\phi_{i}(z;\lambda)\>,$
where the coefficients are given by
$\mathfrak{n}_{i}(\vec{p}_{\perp},\lambda)$ $=$ $i\sqrt{\lambda}\>\alpha_{i}^{(0)}(\vec{p}_{\perp},\lambda)+\sum_{k=1}^{4}\mathfrak{f}_{i}^{(k)}\beta_{k}(\vec{p}_{\perp},\lambda)\>.$ (250)
Comparing (250) with (238), we notice that the difference between $\mathfrak{n}_{i}(\vec{p}_{\perp},\lambda)$ and $\mathfrak{b}_{i}(\vec{p}_{\perp},\lambda)$ is only the sign in front of the first term and therefore the commutation relations among them are computed in a parallel way with the case of (245), i.e.,
$\left[\mathfrak{n}_{i}(\vec{p}_{\perp},\lambda),\mathfrak{n}_{j}^{\>{\dagger}}(\vec{p}_{\perp}^{\>\prime},\lambda^{\>\prime})\right]$ $=$ $0\>,\hskip 42.67912pt(i,j=1,2)\>,$ (251)
$\left[\mathfrak{n}_{i}(\vec{p}_{\perp},\lambda),\mathfrak{b}_{j}^{\>{\dagger}}(\vec{p}_{\perp}^{\>\prime},\lambda^{\>\prime})\right]$ $=$ $\delta^{2}(\vec{p}_{\perp}-\vec{p}_{\perp}^{\>\prime})\delta(\lambda-\lambda^{\>\prime})\times 2\lambda\frac{d\rho_{ij}(\lambda)}{d\lambda}\>.$ (252)
The vanishing commutator (251) indicates that the state $\mathfrak{n}_{i}^{\>{\dagger}}(\vec{p}_{\perp},\lambda)|0\rangle$ is a zero-norm state and is therefore unobservable. The non-trivial commutator (252) implies that two states $\mathfrak{n}_{i}^{\>{\dagger}}(\vec{p}_{\perp},\lambda)|0\rangle$ and $\mathfrak{b}_{i}^{\>{\dagger}}(\vec{p}_{\perp},\lambda)|0\rangle$ have non-vanishing inner product and can communicate with each other. The following four-dimensional commutation relation comes out of (252),
$\thinspace\left[{\cal N}(t,\vec{x}_{\perp},z),\>{\cal B}^{Z}(t^{\>\prime},\vec{x}_{\perp}^{\>\prime},z^{\>\prime})\right]=2i\left\{-\frac{\partial^{2}}{\partial z^{2}}+M_{Z}(z)^{2}\right\}\Delta(t-t^{\>\prime},\vec{x}_{\perp}-\vec{x}_{\perp}^{\>\prime}z,z^{\>\prime};M_{Z}(*))\>.$
The differential operator on the right hand side of (253) is due to the extra $\lambda$ in front of the spectral density function in (252). The equal-time commutation relations
$\left[{\cal N}(t,\vec{x}_{\perp},z),\>{\cal B}^{Z}(t,\vec{x}_{\perp}^{\>\prime},z^{\>\prime})\right]$ $=$ $0\;,$
$\left[{\cal N}(t,\vec{x}_{\perp},z),\>\dot{{\cal B}}^{Z}(t,\vec{x}_{\perp}^{\>\prime},z^{\>\prime})\right]$ $=$ $2i\left\{-\frac{\partial^{2}}{\partial z^{2}}+M_{Z}(z)^{2}\right\}\delta^{2}(\vec{x}_{\perp}-\vec{x}_{\perp}^{\>\prime})\delta(z-z^{\>\prime})\>,$ (254)
are attained immediately from (253).
Incidentally, it is worth remarking that we can derive (253), by starting from the two commutation relations in (254) and by employing the method explained in Appendix C without referring to (252).
6.4 The Slavnov-Taylor identity and the BRST quartet mechanism
In previous sections we have worked out all of the non-vanishing commutation relations between creation and annihilation operators of wave propagation modes associated with the asymptotic fields, ${\cal B}^{Z}(t,\vec{x}_{\perp},z)$, ${\cal N}^{Z}(t,\vec{x}_{\perp},z)$, $c^{Z}(t,\vec{x}_{\perp},z)$, and $\overline{c}^{Z}(t,\vec{x}_{\perp},z)$. Those calculations are suggesting that the states obtained by applying four creation operators
$\mathfrak{b}_{i}^{{\dagger}}(\vec{p}_{\perp},\lambda)\>,\qquad\mathfrak{n}_{i}^{{\dagger}}(\vec{p}_{\perp},\lambda)\>,\qquad\gamma_{i}^{{\dagger}}(\vec{p}_{\perp},\lambda),\qquad\overline{\gamma}_{i}^{{\dagger}}(\vec{p}_{\perp},\lambda)\>,\qquad$ (255)
on a physical state form a BRST quartet. We are now in a position to investigate the decoupling of unphysical states, namely, ghost, anti-ghost, unphysical scalar fields and unphysical polarization states of gauge fields. The point of the decoupling from the physical S-matrix is that, whenever these states appear in the course of time development, they always appear in a particular combination of zero-norm state which is unobservable. The pattern of the combination can be seen by looking at the Slavnov-Taylor type identity slavnov - jctaylor2  .
Let us consider the following identity
$\thinspace\left\{Q_{B},{\rm T}\left({\cal N}(t,\vec{x}_{\perp},z)\overline{c}^{Z}(t^{\>\prime},\vec{x}_{\perp}^{\>\prime},z^{\>\prime})\right)\right\}$ (256)
$=$ ${\rm T}\left(\left[Q_{B},\>{\cal N}(t,\vec{x}_{\perp},z)\right]\overline{c}^{Z}(t^{\>\prime},\vec{x}_{\perp}^{\>\prime},z^{\>\prime})\right)+{\rm T}\left({\cal N}(t,\vec{x}_{\perp},z)\left\{Q_{B},\>\overline{c}^{Z}(t^{\>\prime},\vec{x}_{\perp}^{\>\prime},z^{\>\prime})\right\}\right),$
where $\rm T$}\>" stands for time-ordered product. By sandwiching this identity with physical states, $|{\rm phys}\rangle$ and $\langle{\rm phys}^{\>\prime}|$, and by using the subsidiary condition (87) we arrive at
$0$ $=$ $\langle{\rm phys}^{\>\prime}|{\rm T}\left(\left[Q_{B},\>{\cal N}(t,\vec{x}_{\perp},z)\right]\overline{c}^{Z}(t^{\>\prime},\vec{x}_{\perp}^{\>\prime},z^{\>\prime})\right)|{\rm phys}\rangle$ (257)
$\hskip 28.45274pt+\langle{\rm phys}^{\>\prime}|{\rm T}\left({\cal N}(t,\vec{x}_{\perp},z)\left\{Q_{B},\>\overline{c}^{Z}(t^{\>\prime},\vec{x}_{\perp}^{\>\prime},z^{\>\prime})\right\}\right)|{\rm phys}\rangle\>.$
This identity (257) indicates that the unwanted states created by (255), whenever they appear in graphical calculations, must appear in a particular combination. This can be seen by evaluating the commutator and anti-commutator in (257) explicittly. Before doing so, note that the commutation and anti-commutation relations of $Q_{B}$ with various fields as given in (257) and in (84) are those with Heisenberg fields, while creation and annihilation operators are defined referring to the asymptotic fields. We have to express the relations (84) by using the asymptotic fields.
Let us recall in this connection that Haag haag1 has once argued that Heisenberg operators may be expanded in an infinite series of the asymptotic field products. Using the technique developed in lsz2 , Glaser, Lehmann and Zimmermann (GLZ) glz refined the Haag’s formula further for a wider range of applicability. (See Appendix C in Ref. kugoojima5 for more details on the GLZ formula.) According to the analyses in Refs. kugoojima3 and kugoojima5 , the commutation (or anti-commutation) relations of asymptotic fields with $Q_{B}$ are linear and non-linear interaction terms do not appear. This is plausible since the asymptotic fields are basically free fields. The present case differs from that discussed in kugoojima3 and kugoojima5 in that the mass terms depend on the $z$-coordinate, but still non-linear interaction terms should not appear and the relevant commutator and anti-commutator in (257) then become
$\left\{Q_{B},\overline{c}^{Z}(t,\vec{x}_{\perp},z)\right\}$ $=$ $i\>{\cal B}^{Z}(t,\vec{x}_{\perp},z)\>,$ (258)
$\left[Q_{B},{\cal N}(t,\vec{x}_{\perp},z)\right]$ $=$ $-2i\>\left\{-\frac{\partial^{2}}{\partial z^{2}}+M_{Z}(z)^{2}\right\}c^{Z}(t,\vec{x}_{\perp},z)\>.$ (259)
We note here that the differential operator on the right hand side of (259) is due to the relation (254). This differential operator can be manipulated equivalently in the following way. Namely we note that the mode expansion formula (221) gives
$\thinspace-2i\left\{-\frac{\partial^{2}}{\partial z^{2}}+M_{Z}(z)^{2}\right\}c^{Z}(t,\vec{x}_{\perp},z)$ (260)
$=$ $-2i\int d\lambda\sum_{i=1,2}\int\frac{d^{2}\vec{p}_{\perp}}{\sqrt{(2\pi)^{2}\>2E}}$
$\times\bigg{\{}\lambda\>\gamma_{i}(\vec{p}_{\perp},\lambda)e^{i\vec{p}_{\perp}\cdot\vec{x}_{\perp}-iEt}+\lambda\>\gamma_{i}^{{\dagger}}(\vec{p}_{\perp},\lambda)e^{-i\vec{p}_{\perp}\cdot\vec{x}_{\perp}+iEt}\bigg{\}}\phi_{i}(z;\lambda)\>,$
and this indicates that annihilation and creation operators should be simply replaced by $2i\lambda\gamma_{i}(\vec{p}_{\perp},\lambda)$ and $2i\lambda\gamma_{i}^{{\dagger}}(\vec{p}_{\perp},\lambda)$, respectively, when we use the relation (259).
$\mathfrak{b}_{i}(\vec{p}_{\perp},\lambda)$$\mathfrak{n}_{i}(\vec{p}_{\perp},\lambda)$$\gamma_{i}(\vec{p}_{\perp},\lambda)$$\overline{\gamma}_{i}(\vec{p}_{\perp},\lambda)$$Q_{B}$${\cal N}(t,\vec{x}_{\perp},z)$${\cal B}^{Z}(t,\vec{x}_{\perp},z)$$c^{Z}(t,\vec{x}_{\perp},z)$$\overline{c}^{Z}(t,\vec{x}_{\perp},z)$$Q_{B}$(6.27)(5.66)(6.33)(6.32)(6.26)(5.65)
Figure 2: The relations among the four members of the BRST quartet. The arrows with $Q_{B}$ means that the fields and operators are transformed by the BRST transformation along the arrows. The fields and operators connected by left-right arrows have non-vanishing commutation or anti-commutaion relations.
Thus the identity (257) amounts to saying that, whenever unphysical states appear in the course of time-development, they are always combined together with the following particular relative weight,
$|(\vec{p}_{\perp},\lambda,i),(\vec{p}_{\perp}^{\>\prime},\lambda^{\>\prime},i^{\>\prime})\rangle\equiv\left\{\mathfrak{n}_{i}^{{\dagger}}(\vec{p}_{\perp},\lambda)\mathfrak{b}_{i^{\>\prime}}^{{\dagger}}(\vec{p}_{\perp}^{\>\prime},\lambda^{\>\prime})+2i\lambda\>\gamma_{i}^{\dagger}(\vec{p}_{\perp},\lambda)\overline{\gamma}_{i^{\>\prime}}^{{\dagger}}(\vec{p}^{\>\prime}_{\perp},\lambda^{\>\prime})\right\}|0\rangle\>.$ (261)
The important point here is that (261) is a zero-norm state, i.e.,
$\langle(\vec{q}_{\perp},\lambda_{2},j),(\vec{q}_{\perp}^{\>\prime},\lambda_{2}^{\>\prime},j^{\>\prime})|(\vec{p}_{\perp},\lambda_{1},i),(\vec{p}_{\perp}^{\>\prime},\lambda_{1}^{\>\prime},i^{\>\prime})\rangle=0\>,$ (262)
for arbitrary labels of $(\vec{p}_{\perp},\lambda_{1},i)$, $(\vec{p}_{\perp}^{\>\prime},\lambda_{1}^{\>\prime},i^{\prime})$, $(\vec{q}_{\perp},\lambda_{2},j)$ and $(\vec{q}_{\perp}^{\>\prime},\lambda_{2}^{\>\prime},j^{\prime})$. This can be confirmed by employing the commutation and anti-commutation relations given so far and by computing the following inner products,
$\thinspace\langle 0|\mathfrak{b}_{j^{\>\prime}}(\vec{q}_{\perp}^{\>\prime},\lambda_{2}^{\>\prime})\mathfrak{n}_{j}(\vec{q}_{\perp},\lambda_{2})\mathfrak{n}_{i}^{{\dagger}}(\vec{p}_{\perp},\lambda_{1})\mathfrak{b}_{i^{\>\prime}}^{{\dagger}}(\vec{p}_{\perp}^{\>\prime},\lambda_{1}^{\>\prime})|0\rangle$
$=$ $\delta^{2}(\vec{p}_{\perp}-\vec{q}_{\perp}^{\>\prime})\delta(\lambda_{1}-\lambda_{2}^{\>\prime})\>2\lambda_{1}\frac{d\rho_{ij^{\>\prime}}(\lambda_{1})}{d\lambda_{1}}\cdot\delta^{2}(\vec{p}^{\>\prime}_{\perp}-\vec{q}_{\perp})\delta(\lambda_{1}^{\>\prime}-\lambda_{2})\>2\lambda_{2}\frac{d\rho_{i^{\>\prime}j}(\lambda_{2})}{d\lambda_{2}}\>,$
$\thinspace(+2i\lambda_{1})\cdot(-2i\lambda_{2})\cdot\langle 0|\overline{\gamma}_{j^{\>\prime}}(\vec{q}_{\perp}^{\>\prime},\lambda_{2}^{\>\prime})\gamma_{j}(\vec{q}_{\perp},\lambda_{2})\gamma_{i}^{{\dagger}}(\vec{p}_{\perp},\lambda_{1})\overline{\gamma}_{i^{\>\prime}}^{{\dagger}}(\vec{p}_{\perp}^{\>\prime},\lambda_{1}^{\>\prime})|0\rangle$
$=$ $-4\lambda_{1}\lambda_{2}\cdot\delta^{2}(\vec{p}_{\perp}-\vec{q}_{\perp}^{\>\prime})\delta(\lambda_{1}-\lambda_{2}^{\>\prime})\>\frac{d\rho_{ij^{\>\prime}}(\lambda_{1})}{d\lambda_{1}}\cdot\delta^{2}(\vec{p}^{\>\prime}_{\perp}-\vec{q}_{\perp})\delta(\lambda_{1}^{\>\prime}-\lambda_{2})\>\frac{d\rho_{i^{\>\prime}j}(\lambda_{2})}{d\lambda_{2}}\>.$
The cancellation occurring between (6.4) and (LABEL:eq:gammagammagammagamma) brings us immediately to the zero-norm property (262). The state (261) is thus unobservable. It should also be noted that the state (261) is orthogonal to all of the physical states. The relations among the four members of the BRST quartet are summarized in Figure 2.
7 Physical Polarization States
In the presence of the bubble wall, we are unable to define properly the transverse or longitudinal polarization of gauge field wave-propagation, because of the lack of momentum conservation law in the $z$-direction. However, we are still able to distinguish the physical polarization states from unphysical ones, thanks to the subsidiary condition (87) and the decoupling mechanism of quartet fields from the physical S-matrix explained in the previous section.
First of all it is almost obvious that the states
$\alpha_{i}^{(1)\>{\dagger}}(\vec{p}_{\perp},\lambda)|0\rangle\>,\hskip 28.45274pt\alpha_{i}^{(2)\>{\dagger}}(\vec{p}_{\perp},\lambda)|0\rangle\>,\hskip 28.45274pt(i=1,2)$ (265)
are positive metric states and also satisfy (87) and are therefore physical. In addition to these, we have so far discussed the following states
$\alpha_{i}^{(0)\>{\dagger}}(\vec{p}_{\perp},\lambda)|0\rangle=-\frac{1}{2i\sqrt{\lambda}}\left\{\mathfrak{n}_{i}^{{\dagger}}(\vec{p}_{\perp},\lambda)-\mathfrak{b}_{i}^{{\dagger}}(\vec{p}_{\perp},\lambda)\right\}|0\rangle\>,\hskip 22.76228pt(i=1,2)\>,$ (266)
$\sum_{k=1}^{4}\mathfrak{f}_{i}^{(k)}\beta_{k}^{{\dagger}}(\vec{p}_{\perp},\lambda)|0\rangle=\frac{1}{2}\left\{\mathfrak{n}_{i}^{{\dagger}}(\vec{p}_{\perp},\lambda)+\mathfrak{b}_{i}^{{\dagger}}(\vec{p}_{\perp},\lambda)\right\}|0\rangle\>,\hskip 28.45274pt(i=1,2)\>.$ (267)
We have already seen in Section 6.4 that these states appear in the perturbative calculations only in a particular combination of unobservable zero-norm states together with ghost and anti-ghost states.
Now we would like to argue that there exist other combinations of $\beta_{k}^{{\dagger}}(\vec{p}_{\perp},\lambda)|0\rangle$, i.e.,
$\sum_{k=1}^{4}\mathfrak{h}_{i}^{(k)}\beta_{k}^{{\dagger}}(\vec{p}_{\perp},\lambda)|0\rangle\hskip 28.45274pt(i=1,2)\;,$ (268)
which are orthogonal to (267) and have not yet appeared in the foregoing analyses of the present paper. The coefficients $\mathfrak{h}_{i}^{(k)}(k=1,\cdots,4)$ will be determined soon. In order for the states (268) to be physical, we require the condition
$\left[Q_{B},\>\sum_{k=1}^{4}\mathfrak{h}_{i}^{(k)}\beta_{k}^{{\dagger}}(\vec{p}_{\perp},\lambda)\right]=0,\hskip 28.45274pt(i=1,2)\>.$ (269)
Looking at the expression (83) of $Q_{B}$, it is necessary and sufficient to set
$\sum_{k=1}^{4}\left[\mathfrak{h}_{i}^{(k)}\beta_{k}(\vec{p}_{\perp},\lambda),{\cal B}^{Z}(t,\vec{x}_{\perp},z)\right]=0\;,$ (270)
or equivalently
$\sum_{k=1}^{4}\left[\mathfrak{h}_{i}^{(k)}\beta_{k}(\vec{p}_{\perp},\lambda),\mathfrak{b}_{j}^{{\dagger}}(\vec{p}_{\perp}^{\>\prime},\lambda^{\>\prime})\right]=0$ (271)
for the condition (269) to be satisfied. Putting the definition (238) of $\mathfrak{b}_{j}(\vec{p}_{\perp},\lambda)$ into (271), we get the physical state condition of the state (268)
$\sum_{k=1}^{4}\sum_{l=1}^{4}\mathfrak{h}_{i}^{(k)}\frac{d\widehat{\rho}_{kl}(\lambda)}{d\lambda}\mathfrak{f}_{j}^{(l)}=0\>,$ (272)
which is an algebraic equation to determine the coefficients $\mathfrak{h}^{(k)}_{i}$. In this way we classify the four states $\beta^{{\dagger}}_{k}(\vec{p}_{\perp},\lambda)|0\rangle,\>\>(k=1,\cdots,4)$ into a pair of states given in (267) and another pair of states (268). The latter pair of states contributes to the physical S-matrix, but the former does not. To sum up, physical three pairs of polarization states consist of (265) and (268).
The polarization vectors corresponding to the two states in (265) are of course given by (115) and (124), respectively. On the other hand, the third state (268) is polarized along the $z$-direction and therefore the polarization vector is
[…]
Collecting (115), (124) and (277) altogether, the summation formula over these three physical polarization vectors turns out to be
[…]
Here let us recall that our notations are $g^{\mu\nu}={\rm diag}(1,-1,-1,-1)$ and $\lambda=E^{2}-|\vec{p}_{\perp}|^{2}$. It has been pointed out in farrar that there occurs the rearrangement among transverse and longitudinal polarization states during the passage of the wave through the wall, but our summation formula (282) looks rather simple.
8 Summary
In the present paper we have investigated the standard electroweak gauge field theory in the presence of the bubble wall by taking the $Z^{\mu}$-field sector as the representative case. We included the effect of the position dependent Higgs condensate $v(z)$ into the asymptotic field equations of motion of $Z^{\mu}(t,\vec{x}_{\perp},z)$ and the unphysical scalar field $\chi^{3}(t,\vec{x}_{\perp},z)$. In the $R_{\xi}$-gauge with $\xi=1$, the $Z^{\mu}(t,\vec{x}_{\perp},z)$ asymptotic fields with $\mu=0,1$ and $2$ satisfy the Klein-Gordon type equations of motion (101), but those of $Z^{3}(t,\vec{x}_{\perp},z)$ and $\chi^{3}(t,\vec{x}_{\perp},z)$ are coupled with each other as shown in (142) and (143).
We applied the eigenfunction expansion method developed in Refs. weyl - kodaira2 to the asymptotic field equations and introduced two sets of operators, $\alpha_{i}^{(a)}(\vec{p}_{\perp},\lambda)$ ($i=1,2$ and $a=0,1,2$) and $\beta_{k}(\vec{p}_{\perp},\lambda)$ ($k=1,\cdots,4$). The commutation relations between these operators with their hermitian conjugates are postulated with the help of the spectral functions, as in (5.2.2) and (194). The auxiliary field ${\cal B}^{Z}(t,\vec{x}_{\perp},z)$ that was introduced through the gauge fixing procedure also satisfies the Klein-Gordon type equation (99) in the absence of the tadpole. The annihilation and creation operators, $\mathfrak{b}_{i}$ and $\mathfrak{b}^{{\dagger}}_{i}$, of ${\cal B}^{Z}(t,\vec{x}_{\perp},z)$ are expressed as linear combinations of $\alpha_{i}^{(0)}(\vec{p}_{\perp},\lambda)$ and $\beta_{k}(\vec{p}_{\perp},\lambda)$, and their hermitian conjugates, as shown in (238). We also defined another field ${\cal N}(t,\vec{x}_{\perp},z)$ by (247) and argued that states created by ghost, anti-ghost, ${\cal N}(t,\vec{x}_{\perp},z)$ and the auxiliary ${\cal B}^{Z}(t,\vec{x}_{\perp},z)$ fields constitute the BRST quartet and that they decouple jointly from the physical S-matrix.
Now that the theoretical framework of the standard model gauge theory in the presence of the bubble wall has been put into a good shape, we are now able to launch with renewed interest into reanalyses of plasma interactions with the electroweak bubble wall. We have already witnessed in the last several decades a large variety of publications boedeker1 , boedeker2 - ai3 , dine1 - garcia1 on the coupled dynamics between the bubble wall and the plasma. The microscopic approach based on quantum field theories as given in the present work is appropriate for the ultra-relativistic regime of the wall velocity. It will be interesting to see how the pressure exerted on the wall by plasma particles depends on the Lorentz boost factor, when the effects of multi-soft vector boson emission in the transition radiation are summed up somehow to all orders. A careful look at the infrared region of the phase space is required, and the framework given in this paper will be the most reliable, since the position dependent mass is taken into account without any approximation. For the case of relatively low velocity of the bubble wall, hydrodynamical methods under the local thermal equilibrium condition are often used in literatures. Plasma particles in the distribution function in such approaches are endowed with definite fixed masses and with definite three-momenta. It could be necessary that particles in the distribution function should be labelled with $(\vec{p}_{\perp},\lambda)$ instead of three-momenta, as we did in the present paper, so that the momentum transfer from the plasma particles to the bubble wall should be within the reach in the theoretical framework. We hope we come to these issues in our future publications.
Appendix A The ghost and anti-ghost Lagrangian
For convenience’ sake, we here rewrite the Faddeev-Popov Lagrangian (70) by using the “physical" ghost and anti-ghost fields, i.e., (73) and (74). (See also aoki2 .) The gauge fields are also expressed in terms of the physical combinations (68). The unphysical scalar field, $\chi^{1}$ and $\chi^{2}$, are combined together in the form of (67). The Lagrangian (70) is classified into three terms
${\cal L}_{\rm FP}={\cal L}_{\rm FP}^{(2)}+{\cal L}_{\rm FP}^{(V)}+{\cal L}_{\rm FP}^{(S)}\>,$ (283)
where ${\cal L}_{\rm FP}^{(2)}$ is the quadratic part with respect to the ghost and anti-ghost fields
[…]
$+\partial^{\mu}{\overline{c}}^{Z}\partial_{\mu}c^{Z}-{\overline{c}}^{Z}M_{Z}^{\>2}c^{Z}+\partial^{\mu}{\overline{c}}^{A}\partial_{\mu}c^{A}\>.$
On the other hand, the interaction terms with vector (${\cal L}_{\rm FP}^{(V)}$) and with scalar fields (${\cal L}_{\rm FP}^{(S)}$) are given, respectively, by
${\cal L}_{\rm FP}^{(V)}$ $=$ $\frac{ig}{\sqrt{g^{2}+g^{\>\prime\>2}}}\bigg{[}\partial^{\mu}{\overline{c}}^{(+)}\left\{-c^{(-)}\left(g^{\>\prime}A_{\mu}+gZ_{\mu}\right)+\left(g^{\>\prime}c^{A}+gc^{Z}\right)W_{\mu}^{-}\right\}$ (285)
$+\partial^{\mu}{\overline{c}}^{(-)}\left\{c^{(+)}\left(g^{\>\prime}A_{\mu}+gZ_{\mu}\right)-\left(g^{\>\prime}c^{A}+gc^{Z}\right)W_{\mu}^{+}\right\}$
$+\left(g^{\>\prime}\partial^{\mu}{\overline{c}}^{A}+g\partial^{\mu}{\overline{c}}^{Z}\right)\left(-c^{(+)}W_{\mu}^{-}+c^{(-)}W_{\mu}^{+}\right)\bigg{]}\>,$
${\cal L}_{\rm FP}^{(S)}$ $=$ $\frac{i}{4}g^{2}v\chi^{3}\left({\overline{c}}^{(+)}c^{(-)}-{\overline{c}}^{(-)}c^{(+)}\right)-\frac{1}{4}g^{2}vH\left({\overline{c}}^{(+)}c^{(-)}+{\overline{c}}^{(-)}c^{(+)}\right)$ (286)
$-\frac{1}{4}\left(g^{2}+g^{\>\prime\>2}\right)vH\>{\overline{c}}^{Z}c^{Z}$
$+\frac{igv}{4\sqrt{g^{2}+g^{\>\prime\>2}}}\left(-{\overline{c}}^{(+)}\chi^{-}+{\overline{c}}^{(-)}\chi^{+}\right)\left\{2gg^{\>\prime}c^{A}-\left(g^{\>\prime\>2}-g^{2}\right)c^{Z}\right\}$
$+\frac{igv}{4}\sqrt{g^{2}+g^{\>\prime\>2}}\>{\overline{c}}^{Z}\left(c^{(+)}\chi^{-}-c^{(-)}\chi^{+}\right)\>.$
Appendix B The BRST transformation
The BRST transformation rules of Heisenberg fields given in Section 4.3 are here rewritten by using physical variables, (66), (67), (68), (73) and (74). (The same formulas were previously listed up in aoki2 by using their conventions.) The gauge and scalar field transformation formulas are as follows,
$\delta_{B}W_{\mu}^{\pm}$ $=$ $\delta\lambda\Big{[}\partial_{\mu}c^{(\pm)}\pm\frac{ig}{\sqrt{g^{2}+g^{\>\prime 2}}}\left\{-(g^{\>\prime}c^{A}+gc^{Z})W_{\mu}^{\pm}+(g^{\>\prime}A_{\mu}+gZ_{\mu})c^{(\pm)}\right\}\Big{]}\>,$ (287)
$\delta_{B}Z_{\mu}$ $=$ $\delta\lambda\Big{[}\partial_{\mu}c^{Z}+\frac{ig^{2}}{\sqrt{g^{2}+g^{\>\prime 2}}}\left\{-c^{(+)}W_{\mu}^{-}+c^{(-)}W_{\mu}^{+}\right\}\Big{]}\>,$ (288)
$\delta_{B}A_{\mu}$ $=$ $\delta\lambda\Big{[}\partial_{\mu}c^{A}+\frac{igg^{\>\prime}}{\sqrt{g^{2}+g^{\>\prime 2}}}\left\{-c^{(+)}W_{\mu}^{-}+c^{(-)}W_{\mu}^{+}\right\}\Big{]}\>,$ (289)
[…]
Here $\delta\lambda$ is an anti-commuting c-number. The ghost fields, on the other hand, obey the following transformation rules,
$\delta_{B}c^{(+)}$ $=$ $\delta\lambda\Big{[}-\frac{ig}{\sqrt{g^{2}+g^{\>\prime\>2}}}\left(gc^{Z}+g^{\>\prime}c^{A}\right)c^{(+)}\Big{]}\>,$ (294)
$\delta_{B}c^{(-)}$ $=$ $\delta\lambda\Big{[}\frac{ig}{\sqrt{g^{2}+g^{\>\prime\>2}}}\left(gc^{Z}+g^{\>\prime}c^{A}\right)c^{(-)}\Big{]}\>,$ (295)
$\delta_{B}c^{Z}$ $=$ $\delta\lambda\Big{[}\frac{ig^{2}}{\sqrt{g^{2}+g^{\>\prime\>2}}}c^{(-)}c^{(+)}\Big{]}\>,$ (296)
$\delta_{B}c^{A}$ $=$ $\delta\lambda\Big{[}\frac{igg^{\>\prime}}{\sqrt{g^{2}+g^{\>\prime\>2}}}c^{(-)}c^{(+)}\Big{]}\>,$ (297)
while the rules for the anti-ghost fields are simply given by the auxiliary fields,
$\delta_{B}{\overline{c}}^{(\pm)}$ $=$ ${\color[rgb]{0,0,0}{-}}\delta\lambda B^{\pm}\>,\hskip 28.45274pt\delta_{B}{\overline{c}}^{Z}={\color[rgb]{0,0,0}{-}}\delta\lambda B^{Z}\>,\hskip 28.45274pt\delta_{B}{\overline{c}}^{A}={\color[rgb]{0,0,0}{-}}\delta\lambda B^{A}\>,$ (298)
and those for the auxiliary fields are
$\delta_{B}B^{\pm}=0\>,\hskip 28.45274pt\delta_{B}B^{Z}=0\>,\hskip 28.45274pt\delta_{B}B^{A}=0\>.$ (299)
Appendix C The four-dimensional commutator of ${\cal B}^{Z}(t,\vec{x}_{\perp},z)$
We now prove the four-dimensional commutation relation (218) by using Nakanishi’s technique employed in nakanishinoboru1 and nakanishinoboru2 . We argued in Section 5.1 that ${\cal B}^{Z}(t,\vec{x}_{\perp},z)$ satisfies the Klein-Gordon type equation in the absence of the tadpole, i.e.,
$\big{\{}\square+M_{Z}(z)^{2}\big{\}}{\cal B}^{Z}(t,\vec{x}_{\perp},z)=0\>.$ (300)
Then we can derive the following identity
${\cal B}^{Z}(t,\vec{x}_{\perp},z)$ $=$ $\int d^{2}{\vec{x}}_{\perp}^{\>\prime}\>dz^{\>\prime}\bigg{\{}\frac{\partial}{\partial t^{\>\prime}}\Delta(t-t^{\>\prime},\vec{x}_{\perp}-{\vec{x}}_{\perp}^{\>\prime},z,z^{\>\prime};M_{Z}(*))\>{\cal B}^{Z}(t^{\>\prime},\vec{x}^{\>\prime}_{\perp},z^{\>\prime})$ (301)
$-\Delta(t-t^{\>\prime},\vec{x}_{\perp}-{\vec{x}}_{\perp}^{\>\prime},z,z^{\>\prime};M_{Z}(*))\>\frac{\partial}{\partial t^{\>\prime}}{\cal B}^{Z}(t^{\>\prime},\vec{x}^{\>\prime}_{\perp},z^{\>\prime})\bigg{\}}\>.$
To confirm (301), we first note that the right hand side of (301) is independent of $t^{\>\prime}$, as we see by the following manipulation
$\thinspace\frac{\partial}{\partial t^{\>\prime}}\int d^{2}{\vec{x}}_{\perp}^{\>\prime}\>dz^{\>\prime}\bigg{\{}\frac{\partial}{\partial t^{\>\prime}}\Delta(t-t^{\>\prime},\vec{x}_{\perp}-{\vec{x}}_{\perp}^{\>\prime},z,z^{\>\prime};M_{Z}(*))\>{\cal B}^{Z}(t^{\>\prime},\vec{x}^{\>\prime}_{\perp},z^{\>\prime})$ (302)
$\hskip 42.67912pt-\Delta(t-t^{\>\prime},\vec{x}_{\perp}-{\vec{x}}_{\perp}^{\>\prime},z,z^{\>\prime};M_{Z}(*))\>\frac{\partial}{\partial t^{\>\prime}}{\cal B}^{Z}(t^{\>\prime},\vec{x}^{\>\prime}_{\perp},z^{\>\prime})\bigg{\}}$
$=$ $\int d^{2}{\vec{x}}_{\perp}^{\>\prime}\>dz^{\>\prime}\bigg{\{}\left(\nabla^{\>\prime\>2}-M_{Z}(z^{\>\prime})^{2}\right)\Delta(t-t^{\>\prime},\vec{x}_{\perp}-{\vec{x}}_{\perp}^{\>\prime},z,z^{\>\prime};M_{Z}(*))\>{\cal B}^{Z}(t^{\>\prime},\vec{x}^{\>\prime}_{\perp},z^{\>\prime})$
$\hskip 42.67912pt-\Delta(t-t^{\>\prime},\vec{x}_{\perp}-{\vec{x}}_{\perp}^{\>\prime},z,z^{\>\prime};M_{Z}(*))\>\frac{\partial^{2}}{\partial t^{\>\prime\>2}}{\cal B}^{Z}(t^{\>\prime},\vec{x}^{\>\prime}_{\perp},z^{\>\prime})\bigg{\}}$
$=$ $-\int d^{2}{\vec{x}}_{\perp}^{\>\prime}\>dz^{\>\prime}\Delta(t-t^{\>\prime},\vec{x}_{\perp}-{\vec{x}}_{\perp}^{\>\prime},z,z^{\>\prime};M(*))\>\left\{\square^{\>\prime}+M_{Z}(z^{\>\prime})^{\>2}\right\}{\cal B}^{Z}(t^{\>\prime},\vec{x}^{\>\prime}_{\perp},z^{\>\prime})$
$=$ $0\>.$
Here use has been made of (40), together with the partial integration twice. Since (301) is independent of $t^{\>\prime}$, we are allowed to evaluate the integration in (301) by setting $t=t^{\>\prime}$ and arrive at the left hand side of (301) by using (41) and (42).
Now that (301) has been established, let us compute the commutator by using (301),
$\thinspace\left[{\cal B}^{Z}(t,\vec{x}_{\perp},z),\>{\cal B}^{Z}(t^{\>\prime\prime},\vec{x}_{\perp}^{\>\prime\prime},z^{\>\prime\prime})\right]$ (303)
$=$ $\int d^{2}{\vec{x}}_{\perp}^{\>\prime}\>dz^{\>\prime}\bigg{\{}\frac{\partial}{\partial t^{\>\prime}}\Delta(t-t^{\>\prime},\vec{x}_{\perp}-{\vec{x}}_{\perp}^{\>\prime},z,z^{\>\prime};M_{Z}(*))\>\left[{\cal B}^{Z}(t^{\>\prime},\vec{x}^{\>\prime}_{\perp},z^{\>\prime}),\>{\cal B}^{Z}(t^{\>\prime\prime},\vec{x}_{\perp}^{\>\prime\prime},z^{\>\prime\prime})\right]$
$\qquad-\Delta(t-t^{\>\prime},\vec{x}_{\perp}-{\vec{x}}_{\perp}^{\>\prime},z,z^{\>\prime};M_{Z}(*))\>\frac{\partial}{\partial t^{\>\prime}}\left[{\cal B}^{Z}(t^{\>\prime},\vec{x}^{\>\prime}_{\perp},z^{\>\prime}),\>{\cal B}^{Z}(t^{\>\prime\prime},\vec{x}_{\perp}^{\>\prime\prime},z^{\>\prime\prime})\right]\bigg{\}}\>.$
Since the right hand side of (303) is independent of $t^{\>\prime}$, we can set $t^{\>\prime}=t^{\>\prime\prime}$, use the knowledge of the equal time commutators (217), and are led to conclude that the four-dimensional commutator vanishes, i.e.,
$\left[{\cal B}^{Z}(t,\vec{x}_{\perp},z),\>{\cal B}^{Z}(t^{\>\prime\prime},\vec{x}_{\perp}^{\>\prime\prime},z^{\>\prime\prime})\right]=0\>.$ (304)
This completes the proof of (218).

### v4

@@
 Abstract
-The gauge field theory of the standard electroweak model in the presence of the electroweak bubble wall is investigated with an eye toward its applications to microscopic phenomena, which are supposed to have occurred during the phase transition in the early universe.
+The gauge field theory of the standard electroweak model in the presence of the electroweak bubble wall is investigated with a view to its applications to microscopic phenomena, which are believed to have occurred during the phase transition in the early universe.
 The asymptotic fields are defined anew so that the effects of the position-dependent Higgs condensate are taken into account through the position-dependent $W$ and $Z$ boson masses.
 A novel method of massive gauge field quantization in the $R_{\xi}$-gauge with $\xi=1$ is proposed for the case of the position-dependent masses.
-Our procedure is based on the eigenfunction expansion method associated with second-order differential operators, i.e., a sort of generalized Fourier expansion.The commutation relations of creation and annihilation operators of various wave propagation modes are given in terms of the so-called spectral function.
+Our procedure is based on the eigenfunction expansion method associated with second-order differential operators, i.e., a sort of generalized Fourier expansion.
+The commutation relations of creation and annihilation operators of various wave propagation modes are given in terms of what is known as the spectral function.
 The decoupling of unphysical states from the physical S-matrix is also investigated along the line of Kugo-Ojima’s quartet mechanism on the basis of the BRST symmetry.
@@
 In the presence of the electroweak bubble wall, although we have still the BRST invariance of the Lagrangian, the Lorentz invariance is no more respected and it is far from obvious how the BRST quartet mechanism works.
-The purpose of the present paper is to scrutinize how and to what extent the presence of the electroweak bubble modifies the framework and notions in gauge theories, such as asymptotic fields, field quantization method,  [^1] See Refs.
+The purpose of the present paper is to scrutinize how and to what extent the presence of the electroweak bubble modifies the framework of and notions in gauge theories, such as asymptotic fields, field quantization method,  [^1] See Refs.
 boedeker2 - ai3 for other approaches to field quantization under bubble wall backgrounds.
@@
 $\square\>v+vV^{\>\prime}\left(\frac{v^{2}}{2}\right)=0\>.$ (11)
+By discarding the terms linear in the Higgs field, the separation between background and quantum fluctuations is well defined.
 The trivial constant solutions to (11) are $v=0$ and another one $v=v_{0}\neq 0$ satisfying $V^{\prime}(v_{0}^{\>2}/2)=0$, which correspond to symmetry-restored and symmetry-broken vacuum solutions, respectively.
@@
 In particular we should be concerned with the expansion formulas of interacting Heisenberg field operators with respect to our asymptotic fields haag1 - bogoliubov .
-For now, we have to put our definition of the asymptotic fields on a sound basis by specifying, first of all, a mathematical method of constructing the retarded and advanced Green’s functions.
+For now, we have to put our definition of the asymptotic fields on a sound basis by specifying, first of all, a mathematical method of constructing the retarded and advanced Green’s functions that appear in (LABEL:eq:retardedgreensfunction) and (LABEL:eq:advancedgreensfunction).
 Technical tools for elaborating the Green’s functions are prepared in the course of solving (23) and of getting all the wave propagation modes for given mass function $M_{Z}(z)$.
 Therefore we investigate the solutions to (23) first in Section 3.2 and then will come back later in Section 3.4 to the retarded and advanced Green’s functions.
+A remark to be inserted here for our better understanding is that, in (21) and (22), we take only the temporal infinity ($t\to\pm\infty$) and the spatial coordinates (in particular the $z$-coordinate) are kept fixed.
+This is of course the usual procedure in the LSZ formalism to define the asymptotic fields.
+The overall profile of the Higgs condensate $v(z)$ is unaltered as depicted in Figure 1 and the symmetry-broken and symmetry-restored regions coexist when we solve the wave equation (23).
+For those who would like to compute the friction on the bubble wall in the plasma rest frame, the bubble wall is expanding very rapidly and the Higgs condensate is filling almost all corners of the space in the infinite future.
+Therefore it might be puzzling for them to think of wave propagation modes of (23) which are streching not only in symmetry-broken but also symmetry-restored regions in the $t\to+\infty$ limit.
+In the present paper we are just aiming at the LSZ-like formulation in the presence of the Higgs condensate and the bubble wall rest frame turns out to be the most convenient amenable to the LSZ-like formulation.
+At present, however, we do not have much to say about other Lorentz frames for the use of the LSZ-like formulation.
 3.2 The Klein-Gordon type real scalar field
@@
 should also be mentioned, and these properties enable us to reach (32), (33) and (34) consistently from (37) .
-3.4 The advanced and retarded Green’s functions
+3.4 The retarded and advanced Green’s functions
 We are now well equipped to present the retarded and advanced Green’s functions, which are given by
@@
 $\chi^{\pm}=\frac{\chi^{1}\mp i\chi^{2}}{\sqrt{2}}\>.$ (67)
-The point of rewriting ${\cal L}_{\rm GF}$ as in (62) is that the gauge fields in (63), (64) and (65) are arranged so that they turn out to be $W^{\pm}$-, $Z$- and the electromagnetic ($A_{\mu}$) fields which are related to the original gauge fields $A_{\mu}^{a}$ $(a=1,2,3)$ and $B_{\mu}$ via
-$W_{\mu}^{\pm}\equiv\frac{A_{\mu}^{1}\mp iA_{\mu}^{2}}{\sqrt{2}}\>,\>\>\>Z_{\mu}\equiv\frac{gA_{\mu}^{3}-g^{\>\prime}B_{\mu}}{\sqrt{g^{2}+g^{\>\prime\>2}}}\>,\>\>\>A_{\mu}\equiv\frac{g^{\>\prime}A_{\mu}^{3}+gB_{\mu}}{\sqrt{g^{2}+g^{\>\prime\>2}}}\>,$ (68)
+The point of rewriting ${\cal L}_{\rm GF}$ as in (62) is that the gauge fields in (63), (64) and (65) are arranged so that they turn out to be $W^{\pm}$-, $Z$- and the electromagnetic ($A_{\mu}$)-fields which are related to the original gauge fields $A_{\mu}^{a}$ $(a=1,2,3)$ and $B_{\mu}$ via
+$W_{\mu}^{\pm}\equiv\frac{A_{\mu}^{1}\mp iA_{\mu}^{2}}{\sqrt{2}}\>,\>\>\>Z_{\mu}\equiv\frac{gA_{\mu}^{3}-g^{\>\prime}B_{\mu}}{\sqrt{g^{2}+g^{\>\prime\>2}}}\>,\>\>\>A_{\mu}\equiv\frac{g^{\>\prime}A_{\mu}^{3}+gB_{\mu}}{\sqrt{g^{2}+g^{\>\prime\>2}}}\>.$ (68)
 […]
@@
 4.2 The Faddeev-Popov ghost Lagrangian
-We just follow the standard procedure to write down the Faddeev-Popov Lagrangian ${\cal L}_{FP}$ associated with the gauge fixing (59), that is,
+We just follow the standard procedure to write down the Faddeev-Popov Lagrangian ${\cal L}_{\rm FP}$ associated with the gauge fixing (59), that is,
 ${\cal L}_{\rm FP}$ $=$ $\sum_{a=1}^{3}(\partial^{\mu}{\overline{c}}^{a})\>({\cal D}_{\mu}c)^{a}+(\partial^{\mu}{\overline{c}}^{0})(\partial_{\mu}c^{0})+\frac{1}{4}g^{2}v\sum_{a,b,c=1}^{3}\varepsilon^{abc}\>{\overline{c}}^{a}\chi^{b}c^{c}$ (70)
@@
 $({\cal D}_{\mu}c)^{a}\equiv\partial_{\mu}c^{a}+g\varepsilon^{abc}c^{b}A_{\mu}^{c}\>.$ (71)
-On the ground of argument in Refs.kugoojima1 - kugoojima5 as regard to the (anti-)hermiticity properties of ghost and anti-ghost fields, we take
+On the ground of argument in Refs.kugoojima1 - kugoojima5 in regard to the (anti-)hermiticity properties of ghost and anti-ghost fields, we take
 $c^{a\>\dagger}=c^{a},\qquad\overline{c}^{a\>\dagger}=-\overline{c}^{a},\qquad(a=1,2,3)$
@@
 By construction, the sum of gauge fixing Lagrangian (59) and the Faddeev-Popov Lagrangian (70) is expressed as
-$\delta\lambda\left({\cal L}_{\rm GF}+{\cal L}_{\rm FP}\right)$ $=$ $\delta_{B}\bigg{\{}\sum_{a=1}^{3}\overline{c}^{a}\bigg{(}F^{a}-\frac{1}{2}{\cal B}^{a}\bigg{)}+\overline{c}^{0}\bigg{(}F^{0}-\frac{1}{2}{\cal B}^{0}\bigg{)}\bigg{\}}$ (81)
-$-\delta_{B}\bigg{\{}\partial^{\mu}\bigg{(}\sum_{a=1}^{3}\overline{c}^{a}A^{a}_{\mu}+\overline{c}^{0}B_{\mu}\bigg{)}\bigg{\}}\>,$
+$\delta\lambda\left({\cal L}_{\rm GF}+{\cal L}_{\rm FP}\right)=\delta_{B}\Xi\>,$ (81)
+$\Xi\equiv\sum_{a=1}^{3}\overline{c}^{a}\bigg(F^{a}-\frac{1}{2}{\cal B}^{a}\bigg)+\overline{c}^{0}\bigg(F^{0}-\frac{1}{2}{\cal B}^{0}\bigg)-\partial^{\mu}\bigg(\sum_{a=1}^{3}\overline{c}^{a}A^{a}_{\mu}+\overline{c}^{0}B_{\mu}\bigg)\>,$
 and the invariance of the total Lagrangian density (including the surface terms) under the BRST transformation is obvious by virtue of the nilpotency .
@@
 Note that the fields appearing in $Q_{B}$ and $Q_{c}$ are all interacting Heisenberg fields and the commutation relations of $Q_{B}$ with asymptotic fields have to be discussed later.
+In the previous subsection, we have seen that ${\cal L}_{\rm GF}+{\cal L}_{\rm FP}$ can be expressed as in (81), which is equivalently rewriten as
+$\delta\lambda\left({\cal L}_{\rm GF}+{\cal L}_{\rm FP}\right)=i\left[\delta\lambda Q_{B},\Xi\right]\>.$ (88)
+The subsidiary condition (87) indicates
+$\langle{\rm phys}|\left({\cal L}_{\rm GF}+{\cal L}_{\rm FP}\right)|{\rm phys}^{\prime}\rangle=0\>,$ (89)
+which implies that the effects of adding $\left({\cal L}_{\rm GF}+{\cal L}_{\rm FP}\right)$ to the original symmetric Lagrangian by hand become null in the physical subspace and physical quantities are independent of the choice of the gauge.
+In the present paper, we set $\xi=1$ in the $R_{\xi}$ gauge, but the relations such as (88) and (89) are both valid independently of the way of gauge fixing.
+In fact the proof of the gauge independence of the S-matrix usually starts with (88), but we have to be careful about the gauge dependence of renormalized quantities.
+In the present paper we do not enter the details of the renormalization procedure in the non-trivial $v(z)$ case, and we will discuss the gauge independence of physical quantities after scrutinizing the renormalization in our separate publications.
 5 Field quantization in the presence of the electroweak bubble wall
@@
 $\left(\square+M_{W}^{\>2}\right)W^{\pm}_{\mu}+2M_{W}\frac{\partial_{\mu}v}{v}\chi^{\pm}$ $=$ $0\>,$ (91)
-$\square A_{\mu}$ $=$ $0\>.$ (90)
+$\square A_{\mu}$ $=$ $0\>,$ (92)
 and we notice that unphysical scalar fields appear in (90) and (91) due to $\partial_{\mu}v\neq 0$ .
@@
 In analogy with the scalar case, we postulate the commutation relation
-$\left[\alpha^{(a)}_{i}(\vec{p}_{\perp},\lambda),\;\alpha^{(b)\>{\dagger}}_{j}(\vec{p}_{\perp}^{\>\prime},\lambda^{\>\prime})\right]$ $=$ $-\eta^{ab}\>\delta^{2}(\vec{p}_{\perp}-\vec{p}_{\perp}^{\>\prime})\>\delta(\lambda-\lambda^{\>\prime})\>\frac{d\rho_{ij}(\lambda)}{d\lambda}$
+$\left[\alpha^{(a)}_{i}(\vec{p}_{\perp},\lambda),\;\alpha^{(b)\>{\dagger}}_{j}(\vec{p}_{\perp}^{\>\prime},\lambda^{\>\prime})\right]$ $=$ $-\eta^{ab}\>\delta^{2}(\vec{p}_{\perp}-\vec{p}_{\perp}^{\>\prime})\>\delta(\lambda-\lambda^{\>\prime})\>\frac{d\rho_{ij}(\lambda)}{d\lambda}\>,$
 $(a,b=0,1,2;\>\>\>i,j=1,2)\>,$
@@
 The integration $d\widehat{\rho}_{kl}(\lambda)$ is, precisely speaking, the Stieltjes integral.
-We, however, assume simply that the $\lambda$-spectrum is continuous and therefore we regard (188) as the usual Riemann integration.
+We, however, assume simply that the $\lambda$-spectrum is continuous or equivalently we assume the absence of bound state solutions of the potential $U(z)$ in (157).
+If there would be bound state solutions, they could affect the calculation of the friction on the bubble wall, but for now we leave it for future investigations.
+Under this assumption we regard (190) as the usual Riemann integration.
 We are thus led to the expansion of the pair of the fields in (152) in the following way,
@@
 $=$ $-i\delta^{2}(\vec{x}_{\perp}-\vec{x}_{\perp}^{\>\prime})\delta(z-z^{\>\prime})\>,$
-which is a desirable relation. We also have to confirm that ${\cal B}^{Z}(t,\vec{x}_{\perp},z)$ is a dynamical variable independent of $\chi^{3}(t,\vec{x}_{\perp},z)$ and the commuting property
+which is a desirable relation. We also have to confirm that ${\cal B}^{Z}(t,\vec{x}_{\perp},z)$ is a variable independent of $\chi^{3}(t,\vec{x}_{\perp},z)$ and the commuting property
 $\left[\chi^{3}(t,\vec{x}_{\perp},z),\>\>{\cal B}^{Z}(t,\vec{x}_{\perp}^{\>\prime},z^{\>\prime})\right]=0$ (215)
@@
 The relations among the four members of the BRST quartet are summarized in Figure 2.
+6.5 BRST cohomology classes
+A supplementary remark to be added here is that the zero norm property (264) can also be seen by rewriting (263) in the following way
+$|(\vec{p}_{\perp},\lambda,i),(\vec{p}_{\perp}^{\>\prime},\lambda^{\>\prime},i^{\>\prime})\rangle=-Q_{B}\>\left(\mathfrak{n}_{i}^{{\dagger}}(\vec{p}_{\perp},\lambda)\overline{\gamma}_{i^{\prime}}^{{\dagger}}(\vec{p}_{\perp}^{\>\prime},\lambda^{\prime})|0\rangle\right)\>.$ (267)
+This can be confirmed by noting $Q_{B}|0\rangle=0$ and the following commutation and anti-commutation relations
+$\left[Q_{B},\mathfrak{n}_{i}^{{\dagger}}(\vec{p}_{\perp},\lambda)\right]=-2i\lambda\gamma_{i}^{{\dagger}}(\vec{p}_{\perp},\lambda),\>\>\>\>\left\{Q_{B}\>\overline{\gamma}_{i^{\prime}}^{{\dagger}}(\vec{p}_{\perp}^{\>\prime},\lambda^{\prime})\right\}=-\mathfrak{b}_{i^{\prime}}^{{\dagger}}(\vec{p}_{\perp}^{\prime},\lambda^{\prime}),$ (268)
+which can be deduced, respectively, from (261) and (260). A little more generally one can confirm the identity
+$\thinspace\left\{\mathfrak{n}_{i}^{{\dagger}}(\vec{p}_{\perp},\lambda)\mathfrak{b}_{i^{\>\prime}}^{{\dagger}}(\vec{p}_{\perp}^{\>\prime},\lambda^{\>\prime})+2i\lambda\>\gamma_{i}^{{\dagger}}(\vec{p}_{\perp},\lambda)\overline{\gamma}_{i^{\>\prime}}^{{\dagger}}(\vec{p}^{\>\prime}_{\perp},\lambda^{\>\prime})\right\}|{\rm phys}\rangle$ (269)
+$=$ $-Q_{B}\>\left(\mathfrak{n}_{i}^{{\dagger}}(\vec{p}_{\perp},\lambda)\overline{\gamma}_{i^{\prime}}^{{\dagger}}(\vec{p}_{\perp}^{\>\prime},\lambda^{\prime})|{\rm phys}\rangle\right)$
+for all physical states $|{\rm phys}\rangle$ because of $Q_{B}|{\rm phys}\rangle=0$.
+These are zero norm states and are orthogonal to all of the physical states.
+In gauge theories in general, two physical states, $|{\rm phys}\rangle$ and $|{\rm phys}^{\>\prime}\rangle$, are said to belong to the same BRST cohomology class if the difference between these two states are in the form of $Q_{B}|\psi\rangle$, i.e.,
+$|{\rm phys}\rangle=|{\rm phys}^{\>\prime}\rangle+Q_{B}|\psi\rangle\>.$ (270)
+The states (269) expressed in terms of creation operators all belong to the set of $Q_{B}|\psi\rangle$ states.
 7 Physical Polarization States
@@
 First of all it is almost obvious that the states
-$\alpha_{i}^{(1)\>{\dagger}}(\vec{p}_{\perp},\lambda)|0\rangle\>,\hskip 28.45274pt\alpha_{i}^{(2)\>{\dagger}}(\vec{p}_{\perp},\lambda)|0\rangle\>,\hskip 28.45274pt(i=1,2)$ (265)
-are positive metric states and also satisfy (87) and are therefore physical. In addition to these, we have so far discussed the following states
+$\alpha_{i}^{(1)\>{\dagger}}(\vec{p}_{\perp},\lambda)|0\rangle\>,\hskip 28.45274pt\alpha_{i}^{(2)\>{\dagger}}(\vec{p}_{\perp},\lambda)|0\rangle\>,\hskip 28.45274pt(i=1,2)\>,$ (271)
+are positive norm states and satisfy the physical state condition (87).
+(Note the relations $\left[Q_{B},\alpha_{i}^{(1)\>\dagger}(\vec{p}_{\perp},\lambda)\right]=0$ and $\left[Q_{B},\alpha_{i}^{(2)\>\dagger}(\vec{p}_{\perp},\lambda)\right]=0$.) In addition to (271), we have so far discussed the following states
 $\alpha_{i}^{(0)\>{\dagger}}(\vec{p}_{\perp},\lambda)|0\rangle=-\frac{1}{2i\sqrt{\lambda}}\left\{\mathfrak{n}_{i}^{{\dagger}}(\vec{p}_{\perp},\lambda)-\mathfrak{b}_{i}^{{\dagger}}(\vec{p}_{\perp},\lambda)\right\}|0\rangle\>,\hskip 22.76228pt(i=1,2)\>,$ (272)
@@
 Now we would like to argue that there exist other combinations of $\beta_{k}^{{\dagger}}(\vec{p}_{\perp},\lambda)|0\rangle$, i.e.,
-$\sum_{k=1}^{4}\mathfrak{h}_{i}^{(k)}\beta_{k}^{{\dagger}}(\vec{p}_{\perp},\lambda)|0\rangle\hskip 28.45274pt(i=1,2)\;,$ (268)
+$\sum_{k=1}^{4}\mathfrak{h}_{i}^{(k)}\beta_{k}^{{\dagger}}(\vec{p}_{\perp},\lambda)|0\rangle\>,\hskip 28.45274pt(i=1,2)\;,$ (274)
 which are orthogonal to (273) and have not yet appeared in the foregoing analyses of the present paper.
@@
 which is an algebraic equation to determine the coefficients $\mathfrak{h}^{(k)}_{i}$.
+Eq.
+(278) states that the two sets of states, (273) and (274), are orthogonal to each other.
 In this way we classify the four states $\beta^{{\dagger}}_{k}(\vec{p}_{\perp},\lambda)|0\rangle,\>\>(k=1,\cdots,4)$ into a pair of states given in (273) and another pair of states (274).
@@
 We also defined another field ${\cal N}(t,\vec{x}_{\perp},z)$ by (249) and argued that states created by ghost, anti-ghost, ${\cal N}(t,\vec{x}_{\perp},z)$ and the auxiliary ${\cal B}^{Z}(t,\vec{x}_{\perp},z)$ fields constitute the BRST quartet and that they decouple jointly from the physical S-matrix.
-Now that the theoretical framework of the standard model gauge theory in the presence of the bubble wall has been put into a good shape, we are now able to launch with renewed interest into reanalyses of plasma interactions with the electroweak bubble wall.
-We have already witnessed in the last several decades a large variety of publications boedeker1 , boedeker2 - ai3 , dine1 - garcia1 on the coupled dynamics between the bubble wall and the plasma.
-The microscopic approach based on quantum field theories as given in the present work is appropriate for the ultra-relativistic regime of the wall velocity.
-It will be interesting to see how the pressure exerted on the wall by plasma particles depends on the Lorentz boost factor, when the effects of multi-soft vector boson emission in the transition radiation are summed up somehow to all orders.
+Now that the theoretical framework of the standard model gauge theory in the presence of the bubble wall has been put into a good shape, we are now able to launch with renewed interest into reanalyses of plasma interactions with the electroweak bubble wall that have been dealt with in a large variety of publications boedeker1 , boedeker2 - ai3 , dine1 - doomsday in the last several decades .
+The microscopic approach as given in the present paper is appropriate for studying momentum transfer in $z$-direction from plasma to the bubble wall in a systemmatic way.
+Note that each wave propagation mode of the gauge boson is characterized by $\vec{p}_{\perp}$ and $\lambda$ and there is no conservation law associated with $\lambda$.
+It will be interesting to see how the pressure exerted on the wall by plasma particles depends on the Lorentz boost factor, when the effects of multi-soft vector boson emission in the transition radiation are summed up effectively to all orders.
 A careful look at the infrared region of the phase space is required, and the framework given in this paper will be the most reliable, since the position dependent mass is taken into account without any approximation.
-For the case of relatively low velocity of the bubble wall, hydrodynamical methods under the local thermal equilibrium condition are often used in literatures.
-Plasma particles in the distribution function in such approaches are endowed with definite fixed masses and with definite three-momenta.
-It could be necessary that particles in the distribution function should be labelled with $(\vec{p}_{\perp},\lambda)$ instead of three-momenta, as we did in the present paper, so that the momentum transfer from the plasma particles to the bubble wall should be within the reach in the theoretical framework.
-We hope we come to these issues in our future publications.
+There exists another type of approach based on hydrodynamical methods under the local thermal equilibrium condition.
+This approach is supposed to be useful for the case of relatively low velocity of the bubble wall.
+Plasma particles in the distribution function in such approaches are endowed with definite fixed masses and with definite three-momenta, and the momentum non-conservation in the $z$-direction is introduced in a somewhat ad hoc manner.
+In order to improve such a situation, the field theoretical approach as developed in the present paper should be incorporated in the hydrodynamical approach.
+The quantum version of the Boltzmann equation should be based on the finite temperature Green’s functions, which must be expressed in terms of $\vec{p}_{\perp}$ and $\lambda$ without using such an ill-defined quantity as the $z$-component of momenta.
+Also various techniques, that have been developed in non-equilibrium statistical physics, must be applied extensively to Green’s functions.
+We hope to come to these issues in our future publications.
 Appendix A The ghost and anti-ghost Lagrangian
@@
 Appendix B The BRST transformation
-The BRST transformation rules of Heisenberg fields given in Section 4.3 are here rewritten by using physical variables, (66), (67), (68), (73) and (74).
+The BRST transformation rules of Heisenberg fields given in Section 4.3 are rewritten here by using physical variables, (66), (67), (68), (73) and (74).
 (The same formulas were previously listed up in aoki2 by using their conventions.) The gauge and scalar field transformation formulas are as follows,